### v1

# The Degree of (Extended) Justified Representation and Its Optimization

Abstract
Justified Representation (JR)/Extended Justified Representation (EJR) is a desirable axiom in multiwinner approval voting. In contrast to (E)JR only requires at least one voter to be represented in every cohesive group, we study its optimization version that maximizes the number of represented voters in each group. Given an instance, we say a winning committee provides an (E)JR degree of $c$ if at least $c$ voters in each $\ell$-cohesive group have approved $\ell$ winning candidates. Hence, every (E)JR committee provides the (E)JR degree of at least $1$. Besides proposing this new property, we propose the optimization problem of finding a winning committee that achieves the maximum possible (E)JR degree, called MDJR and MDEJR, corresponding to JR and EJR respectively.
We study the computational complexity and approximability of MDJR of MDEJR. An (E)JR committee, which can be found in polynomial time, straightforwardly gives a $(k/n)$-approximation. On the other hand, we show that it is NP-hard to approximate MDJR and MDEJR to within a factor of $\left(k/n\right)^{1-\epsilon}$, for any $\epsilon>0$, which complements the approximation. Next, we study the fixed-parameter-tractability of this problem. We show that both problems are W[2]-hard if $k$, the size of the winning committee, is specified as the parameter. However, when $c_{\text{max}}$, the maximum value of $c$ such that a committee that provides an (E)JR degree of $c$ exists, is additionally given as a parameter, we show that both MDJR and MDEJR are fixed-parameter-tractable.
1 Introduction
An approval-based committee voting rule (ABC rule) plays a crucial role in collective decision-making by determining a committee from a set of $m$ candidates $C=\{c_{1},\dots,c_{m}\}$, a set of $n$ voters $N=\{1,\dots,n\}$ who each approve a subset $A_{i}$ of $C$, and an integer $k$ representing the desired committee size. Let the list $\mathbf{A}=(A_{1},\dots,A_{n})$ of approval ballots be the ballot profile. Formally, ABC rules take a tuple $(N,C,\mathbf{A},k)$ as input, where $k$ is a positive integer that satisfies $k\leq|C|$, and return one or more size-$k$ subsets $W\subseteq C$, which are called the winning committees. In many concrete voting systems a tiebreaking method is included so that a resolute outcome is guaranteed. In some cases, $N$ and $C$ are omitted from the notation when they are clear from the context. ABC rules are widely applied in various contexts, including the election of representative bodies (such as supervisory boards and trade unions), identifying responses to database queries [12, 34, 19], selecting validators in consensus protocols like blockchain [10], making collective recommendations for groups [24, 25], and facilitating discussions on proposals within liquid democracy [5]. Additionally, as committee elections fall under the domain of participatory budgeting (PB) [9, 18], a strong grasp of ABC rules is indispensable for designing effective PB methods.
The applicability of various ABC rules often depends on the particular context, yet an important requirement one often imposes on an ABC rule is that it can accurately reflect the voters’ preferences, e.g., every large group of voters should justify a seat in the committee. In recent years, a desirable axiom has been proposed and developed, Justified Representation (JR), which requires every group of at least $n/k$ voters that have at least one common candidate should be represented.
Given a profile $\mathbf{A}$ of $n$ approval preferences, a subgroup of voters, denoted by $N^{\prime}\subseteq N$ is called an $\ell$-cohesive group for some $\ell\in\mathbb{N}$, if $|N^{\prime}|\geq\ell\cdot\frac{n}{k}$ and $|\bigcap_{i\in N^{\prime}}A_{i}|\geq\ell$. If $\ell=1$, we say it is a cohesive group for short.
Definition 1 (Justified representation (JR)).
[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies JR for $(\mathbf{A},k)$ if, for every cohesive group (defined right above), there is at least one voter which approves at least one candidate in $W$.
We say that an ABC rule satisfies JR if for each profile $\mathbf{A}$ and committee size $k$, each winning committee provide JR.
Many well-known ABC rules have been shown to satisfy JR, from a simple greedy approval voting rule to a more complex proportional approval voting (PAV) rule. Hence, some stronger JR axioms are proposed to distinguish the existing approval voting rules, such as Extended Justified Representation (EJR). EJR requires that in every $\ell$-cohesive group, at least one member is represented. Hence, EJR implies JR.
Definition 2 (Extended Justified Representation).
[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies EJR for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there are at least one voter which approves at least $\ell$ candidates in $W$.
We say that an ABC rule satisfies EJR if for each profile $\mathbf{A}$ and committee size $k$, each winning committee provide EJR.
While there is a wealth of concepts associated with JR, there are instances where comparing the performance of different winning committees becomes impossible. To illustrate this situation, consider a brief example.
Example 1.
Let $k=1$, $N=\{1,2,3,4\}$, $C=\{c_{1},c_{2},c_{3},c_{4}\}$, $A_{1}=\{c_{1}\}$, and $A_{i}=A_{i-1}\cup\{c_{i}\}$ for every $i=2,3,4$. We are interested in four different committees $W_{i}=\{c_{i}\}$ for every $i\in[4]$.
Since Example 1 is a single winner voting instance, we have $k=\ell=1$. Clearly, $N$ is the only cohesive group in this example, and all four committees in Example 1 satisfy all the above JR axioms: besides the JR and EJR axioms mentioned earlier, one can verify that those committees also satisfy some other JR axioms such as FJR [32], PJR+, and EJR+ [7]. However, $W_{i-1}$ is intuitively better than $W_{i}$ since it satisfies more voters. To distinguish the performance of these committees, besides proposing new JR variants, another approach to addressing this scenario involves leveraging quantitative techniques to model a hierarchy of JR, which has a similar spirit to the work of using quantitative techniques to model proportionality.
Definition 3 (Proportionality degree).
[33] Fix a function $f:\mathbb{N}\rightarrow\mathbb{R}$. An ABC rule has a proportionality degree of $f$ if for each instance $(A,k)$, each winning committee $W$, and each $\ell$-cohesive group $V$, the average number of winners that voters from $V$ approve is at least $f(\ell)$, i.e.,
$\frac{1}{|V|}\sum_{i\in V}|A_{i}\cap W|\geq f(\ell).$
Since $k=1$ in Example 1, it is not hard to see that $W_{1}$, $W_{2}$, $W_{3}$, $W_{4}$ achieve the proportionality degree of $1$, $3/4$, $1/2$, $1/4$, respectively.
Within each cohesive group, the proportionality degree measures the average satisfaction, which may not align with the number of represented voters. Specifically, for a cohesive group where not all voters are represented, optimizing the proportionality degree may further increase the satisfaction of voters who are already represented while leaving some other voters completely unrepresented. The following example demonstrates this.
Example 2.
Let $k=3$, $N=\{1,2,\ldots,9\}$, and $C=\{c_{1},c_{2},\ldots,c_{6}\}$, where
  • $c_{1}$ is approved by voters $1,2,3,4,5$;
  • $c_{2}$ is approved by voters $4,5,6,7,8$;
  • $c_{3}$ is approved by voters $7,8,9,1,2$;
  • each of voters $c_{4},c_{5},c_{6}$ is approved by voters $1,2,4,5,7,8$.
In this example, the winner committee that optimizes the proportionality degree is $\{c_{4},c_{5},c_{6}\}$, where we have $f(1)=2$ and $f(2)=3$, and the value of $f$ is maximized at both $1$ and $2$ (where $f$ is the function defined in Definition 3 but with a fixed instance). In this case, voters $3,6,9$ are completely unsatisfied.
On the other hand, under the winner committee $\{c_{1},c_{2},c_{3}\}$, every voter is represented in both JR and EJR senses. To see this, all voters approve at least one winner, so all voters in each $1$-cohesive group are represented; there is only one $2$-cohesive group $\{1,2,4,5,7,8\}$, and each voter in this cohesive group approves two winners. However, this winner committee is sub-optimal in the proportionality degree measurement: $f(1)=5/3$ and $f(2)=2$.
In the example above, the winner committee $\{c_{4},c_{5},c_{6}\}$ is clearly “unfair” to voters $3,6,9$, and the winner committee $\{c_{1},c_{2},c_{3}\}$ is much more appealing, at least in some applications. Indeed, many social choice scenarios require the committee to represent as many voters as possible (e.g., Monroe’s rule and the social coverage objective) rather than to increase the average satisfaction. Hence, while maximizing average satisfactions of cohesive groups measured by the proportionality degree is a natural goal in some applications, in some other scenarios, increasing the number of represented voters is a more appealing goal. Motivated by these, we seek to define a new quantitative notion that describes the number of represented voters (instead of the average satisfaction) in each cohesive group.
There is also a natural follow-up question: how can we find the committee that satisfies the optimal (maximum) justified representation degree, given an instance $(N,C,\mathbf{A},k)$? In this paper, we propose the degree of (E)JR and study its optimization problem.
1.1 Our New Notions
In this paper, we study the degree of (E)JR and its optimization problem.
Proposal of new notion—(E)JR degree.
Intuitively, given a ballot instance, we say a winning committee provides an (E)JR degree of $c$ if at least $c$ voters in every cohesive group are represented.
Definition 4 (JR Degree).
Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ achieves JR degree $c$ for $(\mathbf{A},k)$ if, for every cohesive group, there are at least $c$ voters each of which approves at least $1$ candidate in $W$.
Definition 5 (EJR degree).
Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ achieves EJR degree $c$ for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there are at least $c$ voters each of which approves at least $\ell$ candidates in $W$.
In this paper, we only consider ballot instances where at least one cohesive group exists, to avoid the uninteresting degenerated case (with no cohesive group) that invalidates the above two definitions.
From the above definitions, a winner committee satisfying JR has a JR degree of at least $1$, and a winner committee satisfying EJR has an EJR degree of at least $1$. In Example 1, $W_{1}$, $W_{2}$, $W_{3}$, $W_{4}$ achieve the (E)JR degree of 4,3,2,1, respectively.
Relationship between JR degree and EJR degree.
In general, given any ballot instance with $n$ voters and a winner committee of size $k$, one can see that both the JR degree and the EJR degree are at most $\lceil\frac{n}{k}\rceil$: if a cohesive group exists, there is always a cohesive group with size exactly $\lceil\frac{n}{k}\rceil$. On the other hand, it is widely known that an EJR committee always exists (and so does a JR committee), and many algorithms are known to find an EJR committee [2, 14] some of which run in polynomial time [3, 32, 7]. Therefore, for any ballot instance, the maximum degrees for both JR and EJR are at least $1$ and at most $\lceil\frac{n}{k}\rceil$.
Since EJR implies JR, it is easy to see from our definitions that, if a winner committee provides an EJR degree of $c$, its JR degree is at least $c$. However, a winner committee may have a higher JR degree than its EJR degree. In addition, the winner committee that maximizes the EJR degree may not be the same as the winner committee that maximizes the JR degree. The following proposition, whose proof is deferred to Appendix B, shows that the difference between the JR degree and the EJR degree can be significant.
Proposition 1.
For any $\gamma>0$, there is a ballot instance with maximum JR degree $c^{\ast}_{\text{JR}}$ and maximum EJR degree $c^{\ast}_{\text{EJR}}$ such that
  1. 1.
    $c^{\ast}_{\text{JR}}/c^{\ast}_{\text{EJR}}>\gamma$,
  2. 2.
    all winner committees with EJR degree $c^{\ast}_{\text{EJR}}$ have JR degrees at most $c^{\ast}_{\text{JR}}/\gamma$, and
  3. 3.
    all winner committees with JR degree $c^{\ast}_{\text{JR}}$ have EJR degrees at most $c^{\ast}_{\text{EJR}}/\gamma$.
While a committee satisfying EJR is generally considered better than one that only satisfies JR, the JR degree and EJR degree are not directly comparable. Intuitively, the EJR degree faces more challenge of satisfying voters in larger cohesive groups, i.e., $\ell$ approved winners needed for representing a voter in a $\ell$-cohesive group, whereas in the JR degree setting, a voter is represented if there is an approved winner. Depending on the scenario, different degree measurements may be more appropriate. Thus, both metrics are valuable and merit further study.
On optimizing (E)JR degree.
The definition of (E)JR degree naturally motivates the following optimization problem, which we define as MDJR/MDEJR.
Definition 6 (Maximum (E)JR degree, MDJR (MDEJR)).
Given an instance $(A,k)$, MDJR (MDEJR) outputs one committee that achieves the maximum (E)JR degree.
“Number” versus “fraction”
We have defined the JR and EJR degrees based on the number of represented voters. Another natural way is to define both notions based on the fraction of represented voters.
When dealing with the JR degree, both definitions are equivalent. To see this, in terms of both numbers and fractions, the least satisfying cohesive group always contains exactly $\lceil n/k\rceil$ voters, as removing a represented voter from a cohesive group of more than $\lceil n/k\rceil$ voters would make both the number and the fraction of represented voters decrease. Given that we are concerning the least satisfying cohesive group (which always has the same size), the “number version” and the “fraction version” of the JR degree are equivalent.
When dealing with the EJR degree, the two definitions are different. The “number version” fits better with the spirit of EJR. In the definition of EJR (see Definition 2), the requirement is that one voter needs to be represented (approve at least $\ell$ winners) in every $\ell$-cohesive group, instead of being that $\ell$ voters need to be represented in every $\ell$-cohesive group. On the other hand, for some $\ell>1$, given an $\ell$-cohesive group with minimum size $\ell\cdot\frac{n}{k}$ and a $1$-cohesive group with minimum size $\frac{n}{k}$, if one voter is represented in the $1$-cohesive group, then $\ell$ voters in the $\ell$-cohesive group need to be represented in order to make the $\ell$-cohesive as “happy” as the $1$-cohesive group in the case the EJR degree is defined in the “fraction version”.
1.2 Our Technical Contributions
In this paper, we focus on the computational complexity and the approximability of the optimization problems MDJR and MDEJR. Our results are listed below.
  • We show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(k/n)^{1-\epsilon}$ for any $\epsilon>0$. This complements the result that an EJR committee (i.e., winner committee with both JR degree and EJR degree at least $1$) always exists and can be computed in polynomial time, which implies a $(k/n)$-approximation for both MDJR and MDEJR.
  • We study the fixed-parameter-tractability of this problem. We show that finding a committee with the maximum achievable (E)JR degree is W[2]-hard if $k$, the size of the winning committee, is specified as the parameter.
  • When the maximum achievable (E)JR degree of an instance is additionally given as a parameter, we show that the problem is fixed-parameter-tractable.
Surprisingly, although Proposition 1 demonstrates that JR degree and EJR degree have different natures, we obtain the same set of results for MDJR and MDEJR.
1.3 Further Related Work
In this subsection, we discuss the related work on justified representation (JR). In addition to the JR axioms previously mentioned, several other JR-related axioms have been proposed and studied. Fernández et al. [17] introduced Proportional JR (PJR), which requires that every $\ell$-cohesive group have some $\ell$ winners represented in the union of their approval sets. PJR is weaker than Extended JR (EJR) but stronger than JR. The authors also proposed Perfect Representation (PR), aiming to represent all voters by some winners, with each winner representing $\frac{n}{k}$ voters. While a PR committee may not always exist, if one does, it can be verified that such a committee guarantees the maximum JR degree of $\frac{n}{k}$.
Peters et al. [32] introduced Fully JR (FJR), which weakens the cohesiveness requirement. It considers groups of $\ell\frac{n}{k}$ voters who share at least $\beta\leq\ell$ candidates in common. A committee satisfies FJR if every $(\ell,\beta)$-weak-cohesive group, where $\ell\in[k]$ and $\beta\in[\ell]$, has at least one member who approves $\beta$ winners. Notably, FJR implies EJR. Brill and Peters [7] proposed PJR+ and EJR+, which focus on ensuring PJR and EJR for groups of $\ell\frac{n}{k}$ voters, where at least one non-elected candidate is approved by all group members, as opposed to considering only $\ell$-cohesive groups. Consequently, EJR+ implies both EJR and PJR+, while PJR+ implies PJR.
Brill et al. [8] studied Individual Representation (IR), which requires that every voter in an $\ell$-cohesive group is represented. An IR committee would achieve the maximum (E)JR degree, although such committees may not always exist. The potential non-existence of IR and PR committees, compared to the guaranteed existence of (E)JR committees, motivates us to explore a quantitative measure bridging (E)JR, PR, and IR.
Moreover, JR has been investigated in other domains, such as fair division [6, 26], participatory budgeting [4, 32], and facility location games [15]. Other properties that assess a committee’s proportionality, such as laminar proportionality and priceability, have also been considered [31].
In our work, we study voting rules that maximize the (E)JR degree, particularly MDJR and MDEJR. In multi-winner voting, there are a variety of voting rules that maximize certain scores, collectively known as Thiele methods [35]. Thiele methods focus on maximizing the sum of voters’ individual satisfaction, where a voter’s satisfaction is determined by the number of approved candidates in the winning committee. While our MDEJR maximization draws inspiration from this concept, MDJR maximization fundamentally differs, as it focuses on maximizing the JR degree within each cohesive group rather than global satisfaction.
Monroe’s rule [28] shares a similar objective with our MDJR approach, as it seeks to maximize the number of voters represented by at least one candidate in the winning committee. This mirrors our goal of maximizing the number of voters represented by at least one winner in every cohesive group. Other notable voting rules, such as Phragmén’s rules [20], have also been studied in the context of multi-winner voting [23].
Finally, several well-known properties in multi-winner approval voting are relevant to our study, including anonymity and neutrality [27, 29, 1], Pareto efficiency [21], monotonicity [16], consistency [22], and strategyproofness [30].
2 Inapproximability of MDJR and MDEJR
We have seen in the previous section that an EJR winner committee always exists and can be computed in polynomial time. Since an EJR winner committee achieves a degree of at least $1$ for both JR and EJR and the maximum possible degree is $n/k$, this implies a $(k/n)$-approximation algorithm for both MDJR and MDEJR problems. To complement this positive result, we will show that both MDJR and MDEJR are NP-hard to approximate to within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
Theorem 1.
It is NP-hard to approximate MDJR within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
Theorem 2.
It is NP-hard to approximate MDEJR within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
We will simultaneously prove the two theorems by constructing a hard ballot instance that is used for proving both theorems. We will make sure the instance we constructed has no $\ell$-cohesive group with $\ell>1$. Notice that the JR degree and the EJR degree for any committee are always the same for instances with only $1$-cohesive groups.
2.1 Proof of Theorem 1 and 2
Before we prove the theorem, we first introduce a NP-hard problem: sparse-SAT problem. One can find variations of SAT problem and a similar argument in [36].
Definition 7 (sparse-SAT).
Given a CNF formula $\phi$ that, for any variables $x$ and $y$, at most one clause contains both $x$ (or $\neg x$) and $y$ (or $\neg y$), decide if there is a value assignment to the variables to make $\phi$ true.
To see the NP-hardness, it can be reduced from the SAT problem. Start with any SAT instance. Without loss of generality, suppose each variable, $x$ or $\neg x$, appears in each clause at most once. For each variable $x$ that there are more than two clauses containing $x$ or $\neg x$ perform the following procedure: Suppose $x$ appears in $k$ clauses. Create $k$ new variables $x_{1},\dots,x_{k}$ and replace the $i$th occurrence of $x$ with $x_{i}$ (and $\neg x$ is replaced by $\neg x_{i}$, respectively), $i=1,\dots,k$. Append the clause $(x_{i}\vee\neg x_{i+1})$ for $i=1,\dots,k-1$ and the clause $(x_{k}\vee\neg x_{1})$. Note that, in the new instance, variable $x_{i}$ and $y_{j}$ appear in a clause only when the $i$th occurrence of $x$ and the $j$th occurrence of $y$ in $\phi$ are in the same clause, so the new instance satisfies the limitation of sparse-SAT problem.
In the new instance, the clause $(x_{i}\vee\neg x_{i+1})$ implies that if $x_{i}$ is false, $x_{i+1}$ must be false as well. The cyclic structure of the clauses therefore forces the $x_{i}$ to be either all true or all false, so the new instance is satisfiable if the original one is. Moreover, the transformation requires polynomial time.
Now we are ready to prove our theorems.
Proof of Theorem 1 and 2.
We reduce from sparse-SAT problem. Given any sparse-SAT instance $\phi$, suppose there are $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$). We consider a ABC voting instance with $13\bar{m}+\bar{n}+$ 1 candidates, $c_{1},\dots,c_{2\bar{m}}$, $s_{1},\dots,s_{\bar{n}}$, $t_{1},\dots,t_{\bar{m}}$, $d$ and $\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}$ voters. We want to select a committee of size $\bar{m}+1$. Hence, we care about the cohesive group of size $\frac{\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}}{\bar{m}+1}=\bar{n}+\bar{m}$. First, for each variable $x_{j}$ and its complement $\neg x_{j}$, we create two corresponding candidates $c_{2j-1}$ and $c_{2j}$ and a group $T_{j}$ of $\bar{m}$ voters who approve $c_{2j-1}$ and $c_{2j}$. For the $i$th clause, we create a group $S_{i}$ of $\bar{m}$ voters. All voters in group $S_{i}$ approve $c_{2j-1}$ if $x_{j}$ occurs in the $i$th clause and $c_{2j}$ if $\neg x_{j}$ occurs in the $i$th clause. All voters in group $S_{i}$ approve $s_{i}$ and voters in group $T_{i}$ approve $t_{i}$ additionally. We create a set $D$ of $\bar{n}$ voters who approve $s_{1},\dots,s_{\bar{n}},t_{1},\dots,t_{\bar{m}},d$. Hence, for each $i\in[\bar{n}]$, $S_{i}\cup D$ forms a $1$-cohesive group and for each $j\in[\bar{m}]$, $T_{j}\cup D$ forms a $1$-cohesive group. Moreover, we create a set $D^{+}$ of $\bar{m}$ voters who approve $d$. Hence, $D\cup D^{+}$ forms a $1$-cohesive group.
Notably, there is no $2$-cohesive group. First, candidate $s_{i}$ for $i\in[\bar{n}]$ or $t_{j}$ for $j\in[\bar{m}]$ or $d$ has only $\bar{n}+\bar{m}$ voters approving them as constructed above. For candidates $c_{1},\dots,c_{2\bar{m}}$, we will show that no $2(\bar{n}+\bar{m})$ voters have two common approved candidates. Since in a sparse-SAT instance, for any two variables $x_{i}$ and $x_{j}$, at most one clause contains them or their complement, for any two candidates $c_{2i}$ or $c_{2i-1}$ and $c_{2j}$ or $c_{2j-1}$, at most one group $S_{k}$ of $\bar{m}$ voters approves them simultaneously. Moreover, only $\bar{m}$ voters in group $T_{i}$ approve both $c_{2i-1}$ and $c_{2i}$. So there is no $2$-cohesive group or $\ell$-cohesive group for $l>2$. In this case, the EJR degree equals to JR degree.
If there is a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of $\bar{n}+\bar{m}$. To see this, we elect $c_{2j-1}$ as a winner if $x_{j}$ assigned true or we elect $c_{2j}$ for all $j\in[\bar{m}]$. Then we elect $d$ as the winner. We can verify that all voters in $S_{i}$ and $T_{j}$ are covered for every $i\in[\bar{n}],j\in[\bar{m}]$. In addition, $d$ covers all voters in $D$ and $D^{+}$. Hence, all voters are covered and the JR degree equals the size of $1$-cohesive group, $\bar{n}+\bar{m}$, and so does the EJR degree.
If there does not exist a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of at most $\bar{n}$. To see this, we prove it by contradiction. Assume that MDEJR can achieve the EJR degree larger than $\bar{n}$. If $d$ is not elected, no voter in $D^{+}$ can be covered since they only approve $d$. Hence, for cohesive group $D\cup D^{+}$, there are at most $\bar{n}$ voters can be covered, which makes a contradiction. If $d$ is elected, there will be $\bar{m}$ vacant seats. If voters in $T_{j}$ are not covered for $j\in[\bar{m}]$, there are at most $\bar{n}$ voters covered in cohesive group $D\cup T_{j}$, a contradiction, so all voters in group $T_{j}$ must be covered for all $j\in[\bar{m}]$, indicating that exact one of $3$ candidates $\{c_{2j-1},c_{2j},t_{j}\}$ is elected. Next, we will show that there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered. To see this, suppose there is a committee $W^{\prime}$ of size $\bar{m}$ which covers all groups $S_{1},\dots,S_{\bar{n}}$. For every $t_{j}\in W^{\prime}$, we can find $c_{2j-1}$ or $c_{2j}$ to replace $t_{j}$ since $t_{j}$ only covers group $T_{j}$. Hence, we can find one candidate in $\{c_{2i-1},c_{2i}\}$ for all $i\in[\bar{m}]$ to cover all groups, implying that there exists a value assignment to the variables ($x_{i}$ is assigned true if $c_{2i-1}$ is elected or false otherwise) to make $\phi$ true, leading to a contradiction. Therefore, there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered. Without loss of generality, we assume that it is $S_{i}$. Then, for cohesive group $S_{i}\cup D$, there are at most $\bar{n}$ voters that can be covered, which makes a contradiction.
From the proof of NP-hardness, we further have MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}}{\bar{n}+\bar{m}}$ (Note that $\bar{n}$ and $\bar{m}$ are the parameters of the sparse-SAT problem rather than ABC voting, we will reformulate it later). Now, instead of reducing from sparse-SAT problems directly, we add an intermediate reduction. Given a sparse-SAT instance $\phi$ with $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$), there is a corresponding ‘overstaffing’ sparse-SAT instance $\phi^{\prime}$ with $\bar{n}^{\prime}=\bar{n}+1$ clauses and $\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ different variables such that a new clause is added with a number of $(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ new variables like $(y_{1}\vee y_{2}\vee\dots)$. Obviously, the ‘overstaffing’ sparse-SAT problem is also an NP-hard problem.
Now we reduce our problem from ‘overstaffing’ sparse-SAT problem in the same way that we reduce from sparse-SAT problem. Similarly, MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}$ where $\bar{n}^{\prime}=\bar{n}+1,\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$. Recall the construction, the number of voters and the committee size are $n_{\text{voting}}=\bar{n}^{\prime}\bar{m}^{\prime}+\bar{m}^{\prime 2}+\bar{n}^{\prime}+\bar{m}^{\prime}$ and $k_{\text{voting}}=\bar{m}^{\prime}+1$, respectively, which can be reformulated as $\frac{k_{\text{voting}}}{n_{\text{voting}}}=\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}$. Since
$(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}=(\bar{n}^{\prime}+\bar{m}+(\bar{n}^{\prime}+\bar{m})^{\lceil\frac{1}{\epsilon}\rceil})^{\epsilon}>\bar{n}^{\prime}+\bar{m}>\bar{n}^{\prime}.$
Finally, MDJR cannot be approximated in polynomial time to within a factor of
$\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}\leq\frac{(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}}{\bar{n}^{\prime}+\bar{m}^{\prime}}=\left(\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}\right)^{1-\epsilon}\!\!\!=\!\left(\frac{k_{\text{voting}}}{n_{\text{voting}}}\right)^{1-\epsilon},$
where $\epsilon>0$. ∎
3 Parameterized Complexity of MDJR/MDEJR
The parameterized approach is often used to address problems that are hard to solve in their general form but become more tractable or have improved algorithms when considering specific parameter values. In most scenarios, the committee size $k$ is much smaller than the number of voters. Hence, would it be helpful if we fixed the parameter $k$?
In Appendix A, we provide a brief introduction to the parameterized complexity theory which includes basic definitions that are necessary for the results in the remaining part of our paper. The readers familiar with parameterized complexity can move on to Section 3.1.
3.1 W[2]-Hardness with Parameter $k$.
We first show that both MDJR and MDEJR are intractable when the committee size $k$ is specified as a parameter.
Theorem 3.
MDJR is W[2]-hard parameterized by $k$.
Proof.
We present a reduction from the set cover problem, a canonical W[2]-complete problem [11]. Given a set cover instance $(\mathcal{U}=\{1,\ldots,n\},\mathcal{S}=\{S_{1},\ldots,S_{m}\},k)$, we construct an MDJR instance $(N,C,\mathbf{A},k^{\prime})$ as follows.
The set of voters $N$ is given by $N=N_{1}\cup N_{2}$ where $N_{1}=\{1,\ldots,n\}$ corresponds to $\mathcal{U}$ in the set cover instance and $N_{2}=\{n+1,\ldots,2n\}$ is the set of $n$ additional voters. The set of candidates $C=\{c_{1},\ldots,c_{m}\}$ corresponds to $\mathcal{S}$. The profile $\mathbf{A}$ is defined as follows:
  • for voters in $N_{1}$, a voter $i$ approves a candidate $c_{j}$ if and only if $i\in S_{j}$ in the set cover instance;
  • for every voter in $N_{2}$, they approve all candidates.
Finally, the committee size $k^{\prime}$ is set to $k^{\prime}=k$. We assume $k\geq 2$ without loss of generality.
If the set cover instance is a yes instance, by selecting the $k$ candidates corresponding to the $k$ subsets that cover $\mathcal{U}$, every voter approves at least one candidate in the winner committee. The JR degree reaches the maximum possible value $|N|/k^{\prime}$.
If the set cover instance is a no instance, for every winner committee with $k^{\prime}=k$ candidates, there exists a voter in $N_{1}$ that approves no candidate in the winner committee. On the other hand, every voter in $N_{1}$ is in at least one cohesive group, as this voter and the $n$ voters in $N_{2}$, with a total of $n+1>|N|/k^{\prime}$ voters, approve one common candidate. This implies the JR degree cannot reach the maximum possible value $|N|/k^{\prime}$. ∎
The hardness of MDEJR is proved differently. In the previous section, we prove the hardness of MDEJR in the same way as MDJR by making sure only $1$-cohesive groups exist. However, we fail to make this technique work for proving the following theorem. Instead, the hard instance constructed here contains $\ell$-cohesive groups for $\ell>1$.
Theorem 4.
MDEJR is W[2]-hard parameterized by $k$.
Proof.
We again present a reduction from the W[2]-complete problem, the set cover problem. We make the following assumptions on the set cover instance $(\bar{U}=\{1,\ldots,\bar{n}\},\bar{S}={S_{1},\ldots,S_{\bar{m}}},\bar{k})$ without loss of generality.
  • $\bar{k}<\bar{m}$. Otherwise, the set cover instance is trivial.
  • $\bar{n}$ is a multiple of $9$. To achieve this, we can just add dummy elements that are covered by all the subsets.
  • For any two subsets $S_{i},S_{j}\in\bar{S}$, we have $|S_{i}\cap S_{j}|\leq\frac{1}{75}|\bar{U}|=\bar{n}$. To achieve this, for the original instance, we can create $74\bar{n}$ additional elements, create a subset that contains these elements, and increase $\bar{k}$ by $1$.
Given a set cover instance $(\bar{U}=\{1,\ldots,\bar{n}\},\bar{S}={S_{1},\ldots,S_{\bar{m}}},\bar{k})$ satisfying the above three assumptions, we construct an MDEJR instance as follows. Let $\bar{n}^{\prime}=\frac{4}{3}\bar{n}$, and notice that $\bar{n}^{\prime}$ is a multiple of $3$.
The MDEJR instance has $n=\bar{n}^{\prime}\cdot(\bar{k}+3)$ voters which form four groups: $U$, $U^{\prime}$, $V$, and $W$. It has $m=\bar{m}+4$ candidates, $c_{1},\ldots,c_{\bar{m}},$ $c^{\ast},d_{1},d_{2},d_{3}$. The size of the winner committee $k$ is set to $k=\bar{k}+3$. Voter group $U$ and candidates $c_{1},\ldots,c_{\bar{m}}$ correspond to the set cover instance, where voter $i$ approves candidate $c_{j}$ if and only if $i\in S_{j}$ in the set cover instance. In addition, all voters in $U$ approve $c^{\ast}$. Voter group $U^{\prime}$ consists of $\bar{n}^{\prime}-\bar{n}=\frac{1}{3}\bar{n}$ voters, who only approve $c^{\ast}$. Voter group $V$ consists of $(k-3)\cdot\frac{n}{k}$ voters, and each voter in $V$ approves $\{c_{1},\ldots,c_{\bar{m}}\}$. Finally, voters in $W$ are partitioned into $6$ groups $W_{1},\ldots,W_{6}$ such that each group contains $\frac{n}{3k}$ voters. Candidate $d_{1}$ is approved by voters in $W_{1}\cup W_{2}\cup W_{3}$, candidate $d_{2}$ is approved by voters in $W_{3}\cup W_{4}\cup W_{5}$, and candidate $d_{3}$ is approved by voters in $W_{5}\cup W_{6}\cup W_{1}$. Since we have $|U|=\bar{n}=\frac{3n}{4k}$ and $|U^{\prime}|=\frac{1}{3}\bar{n}=\frac{n}{4k}$, the total number of voters sums up to $n$:
$|U|+|U^{\prime}|+|V|+|W|=\frac{3n}{4k}+\frac{n}{4k}+(k-3)\frac{n}{k}+6\cdot\frac{n}{3k}=n.$
Since we have seen that $\bar{n}$ is a multiple of $3$, $\frac{n}{3k}$ is an integer.
We list the following two key observations:
  1. 1.
    To make EJR degree positive, we need to select at least $\bar{k}$ candidates from $\{c_{1},\ldots,c_{\bar{m}}\}$. To see this, voters in $V$ approve $\bar{m}\geq\bar{k}=k-3$ common candidates, so $V$ is a $(k-3)$-cohesive group since $|V|=(k-3)\cdot\frac{n}{k}$. If less than $\bar{k}=k-3$ candidates are selected from $\{c_{1},\ldots,c_{\bar{m}}\}$, none of the voters in $V$ is represented, and the EJR degree becomes 00.
  2. 2.
    If less than $3$ candidates are selected from $\{d_{1},d_{2},d_{3}\}$, the EJR degree is at most $\frac{2n}{3k}$. To see this, the three sets of voters, $W_{1}\cup W_{2}\cup W_{3}$, $W_{3}\cup W_{4}\cup W_{5}$, and $W_{5}\cup W_{6}\cup W_{1}$, are $1$-cohesive groups with size exactly $\frac{n}{k}$. Since voters in $W_{2}$, $W_{4}$, and $W_{6}$ have only one approved candidate $d_{1}$, $d_{2}$, and $d_{3}$ respectively, if one of $\{d_{1},d_{2},d_{3}\}$ is not selected, voters from one of $W_{2}$, $W_{4}$, and $W_{6}$ are not represented, which implies the EJR degree is at most $2\cdot\frac{n}{3k}$.
By the observations above, to have a winner committee with the EJR degree at least $\frac{2n}{3k}$, we need to select exactly $\bar{k}$ candidates from $c_{1},\ldots,c_{\bar{m}}$ and all the $3$ candidates $d_{1},d_{2},d_{3}$. In particular, we have no chance to select $c^{\ast}$, which induces the $1$-cohesive group $U\cup U^{\prime}$ with $|U\cup U^{\prime}|=\frac{n}{k}$.
If the set cover instance is a yes instance, we will show that we can achieve an EJR degree of at least $\frac{3n}{4k}$. We select those $\bar{k}$ candidates in $\{c_{1},\ldots,c_{\bar{m}}\}$ that correspond to the set cover solution, and we additionally select $d_{1},d_{2}$, and $d_{3}$. For all the three $1$-cohesive groups within $W$, all voters are represented, and the EJR degree here is $\frac{n}{k}>\frac{3n}{4k}$. For the $1$-cohesive group $U\cap U^{\prime}$, all voters in $U$ are represented (since the set cover instance is a yes instance), and the EJR degree here is exactly $|U|=\bar{n}=\frac{3}{4}\bar{n}^{\prime}=\frac{3n}{4k}$. Now we reason about those $\ell$-cohesive groups with $\ell>1$. Notice that, for at least $2\cdot\frac{n}{k}$ voters to approve at least $2$ common candidates, the $2$ common candidates can only come from $\{c_{1},\ldots,c_{\bar{m}}\}$ (since each of $c^{\ast},d_{1},d_{2},d_{3}$ is approved by only $n/k$ voters). For any $c_{i},c_{j}\in\{c_{1},\ldots,c_{\bar{m}}\}$, by our third assumption on the set cover instance, the voters that approve both $c_{i}$ and $c_{j}$ are those in $V$ together with at most $\frac{1}{75}\bar{n}$ voters in $U$. Note also that all voters in $V$ approve $\bar{k}=k-3$ candidates in the winner committee and $k-3$ is the maximum value of $\ell$ for the existence of $\ell$-cohesive groups. Given any $\ell$-cohesive group with $\ell>1$, at most $\frac{1}{75}\bar{n}$ voters are not represented. Thus, the EJR degree for those $\ell$-cohesive groups with $\ell>1$ is at least
$\frac{n}{k}-\frac{1}{75}\cdot\bar{n}=\frac{n}{k}-\frac{1}{75}\cdot\frac{3n}{4k}=\frac{99n}{100k}>\frac{3n}{4k}.$
In conclusion, we have shown that the EJR degree for the committee constructed above is $\frac{3n}{4k}$.
If the set cover instance is a no instance, we will show that the EJR degree for any committee is strictly less than $\frac{3n}{4k}$. Suppose for the sake of contradiction that this is not the case, and we have a winner committee with an EJR degree of at least $\frac{3n}{4k}$. Since $\frac{3n}{4k}>\frac{2n}{3k}$, our two observations implies that we have to select $\bar{k}$ candidates from $c_{1},\ldots,c_{\bar{m}}$ and we have to select all of $d_{1},d_{2},d_{3}$. Since the set cover instance is a no instance, at least one voter in $U$ has no approved candidate in the winner committee. Moreover, since $c^{\ast}$ is not selected, all voters in $U^{\prime}$ do not have an approved winner candidate. Therefore, in the $1$-cohesive group $U\cup U^{\prime}$ of size $n/k$, at least $11+|U^{\prime}|=1+\frac{1}{3}\bar{n}=\frac{n}{4k}+$ 1 voters are not represented. As a result, the EJR degree is at most $\frac{3n}{4k}-1$, which contradicts our assumption that the EJR degree is at least $\frac{3n}{4k}$. ∎
3.2 Fixed-Parameter-Tractability with Parameters $k$ and Maximum (E)JR Degree
We have seen that both MDJR and MDEJR are still computationally hard even parameterized by $k$. Thus, to make the problems tractable, different choices of the parameters or additional parameters are needed.
If we choose the number of candidates $m$ as the parameter, it is easy to verify that both MDJR and MDEJR are fixed-parameter-tractable. To see this, we can enumerate all the $\binom{m}{k}$ committees. For each committee, we can compute the (E)JR degree in $O(2^{m}m^{2}n)$ time [3]. At last, we select the committee that achieves the maximum (E)JR degree.
Another natural choice for the parameter is the maximum achievable (E)JR degree. Fortunately, both MDJR and MDEJR become tractable if parameterized by both $k$ and the maximum (E)JR degree. In the next two sections, we use $c_{\text{max}}$ to denote the maximum (E)JR degree.
3.2.1 Algorithm for MDJR
Our starting point is the algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2], which always outputs a JR committee.
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow\emptyset$
2for $j\in[k]$ do
3       Let $c$ be the candidate approved by the maximum number of voters in $N$
4      $W\leftarrow W\cup c$
5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
return $W$
Algorithm 1 Greedy Approval Voting (GreedyAV)
Next, we show the following property for GreedyAV which is the key for our algorithm. It states that the algorithm GreedyAV also gives us the optimal JR degree if $n$ is large enough.
Proposition 2.
Given any instance with the maximum achievable JR degree of $c_{\emph{max}}$ and $n>k^{2}(c_{\emph{max}}-1)$, GreedyAV will output the committee achieving JR degree $c_{\emph{max}}$.
Proof.
We prove it by contradiction. Suppose the committee output by GreedyAV provides the JR degree less than $c_{\text{max}}$. Hence, there will be at least $(\frac{n}{k}-c_{\text{max}}+1)$ voters that approve a non-elected candidate. From the definition of GreedyAV, the coverage of voters is at least $(\frac{n}{k}-c_{\text{max}}+1)$ in each iteration, and at least $\frac{n}{k}$ in the first iteration. Therefore, the total number of voters is at least
$\frac{n}{k}+(k-1)\left(\frac{n}{k}-c_{\text{max}}+1\right)+\left(\frac{n}{k}-c_{\text{max}}+1\right)$
$=n+\frac{n}{k}-k(c_{\text{max}}-1)>n,$
which leads to a contradiction in the number of voters. ∎
Given any instance with $n>k^{2}(c_{\text{max}}-1)$, GreedyAV can achieve the maximum JR degree. Hence, the remaining case is $n\leq k^{2}(c_{\text{max}}-1)$. Given any instance $(N,C,\mathbf{A},k)$, each candidate $c_{i}\in C$ can be seen as a subset of $N$, including all voters that approve $c_{i}$. Hence, there are at most $2^{n}$ different types of candidates, implying that every instance corresponds to an equivalent instance with $m\leq 2^{n}$. Therefore, we can decide whether there exists a committee providing JR degree of $c$ by enumerating all the committees when $n\leq k^{2}(c-1)$, which can be computed with running time
$\binom{m}{k}\leq\binom{2^{n}}{k}\leq\binom{2^{k^{2}(c-1)}}{k}=f(k,c).$
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow\text{GreedyAV}(\mathcal{I})$
2for $c:\lceil\frac{n}{k^{2}}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
3       Enumerate all those $\binom{m}{k}$ possible committees to see if there is a committee $W^{*}$ achieving JR degree $c$;
4       if $W^{*}$ exists then
5             $W\leftarrow W^{*}$
6      else
7             return $W$
8
Algorithm 2 MDJR Voting Rule
Our algorithm is described in Algorithm 2. We prove that Algorithm 2 outputs the committee that achieves the maximum JR degree and runs in time $f(k,c_{\text{max}})\cdot\text{poly}(m,n)$.
Running Time.
GreedyAV can be computed in polynomial time. It is easy to see that the number of for loop is bounded by $O(\frac{n}{k})$. Then, we only need to justify the running time of deciding the condition of if branch. Since $c>\frac{n}{k^{2}}$, we have $n<k^{2}c$. As we mentioned before, we can just enumerate all committees in running time $\binom{2^{k^{2}c}}{k}$ and decide the JR degree for each committee in polynomial time.
Correctness.
If MDJR voting rule executes else branch in the first iteration of for loop, there does not exist a committee with JR degree $c$ for $c>\frac{n}{k^{2}}$. Hence, we have $c_{\text{max}}\leq\frac{n}{k^{2}}$. From Proposition 2, we know that GreedyAV achieves JR degree $c_{\text{max}}$. For the other cases, suppose that the else branch is executed when $c=c_{\text{end}}$. Then there does not exist a committee with JR degree $c_{\text{end}}$. The latest $W$ is the committee that provides JR degree $(c_{\text{end}}-1)$, which achieves the maximum JR degree.
3.2.2 Algorithm for MDEJR
Similar to MDJR, we first consider an EJR voting rule, proportional approval voting (PAV) [2], which outputs the committee that maximizes the PAV-score, where the PAV-score of a committee $W\subseteq C$ is defined as
$s_{\text{PAV}}(W)=\sum_{i=1}^{n}\sum_{j=1}^{|A_{i}\cap W|}\frac{1}{j}.$
In the PAV-score, each voter’s “utility” is defined by the harmonic progression $H[t]$ for $t$ approved winning candidates and the PAV-score can then be understood as the social welfare. One may wonder whether PAV can provide the maximum EJR degree. We find that PAV fails to achieve the maximum EJR degree in some instances. A counterexample can be found in Appendix C.
In addition, PAV cannot be computed in polynomial time. Aziz et al. [3] showed that a local search alternative algorithm for PAV can both satisfy EJR and be computed in polynomial time. The algorithm is described in Algorithm 3. Starting from an arbitrary winning committee, the algorithm considers all possible single-candidate-replacements that increase the PAV score by at least $\lambda$ (where $\lambda$ is a parameter of the algorithm). For each pair of candidate $c^{+}$ and $c^{-}$ with $c^{+}\notin W$ and $c^{-}\in W$, if we swap $c^{+}$ and $c^{-}$, i.e. to remove $c^{-}$ from the committee and select $c^{+}$ instead, the score is increased by $\Delta(W,c^{+},c^{-})=s_{\text{PAV}}(W\setminus\{c^{-}\}\cup\{c^{+}\})-s_{\text{PAV}}(W).$
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow k$ arbitrary candidates from $C$
2while there exist $c^{+}\notin W$ and $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$ do
3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
return $W$
Algorithm 3 $\lambda$-LS-PAV
Next, we prove the following observation in a similar spirit to Proposition 2, which shows that the local search variant of PAV can achieve maximum EJR degree if $n$ is sufficiently large and $\lambda$ is sufficiently small.
Proposition 3.
Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$. For all $\lambda<\frac{n}{k^{2}}$ satisfying $n\geq k(k+1)(c_{\emph{max}}-1)+\lambda k^{2}$, $\lambda$-LS-PAV will output the committee achieving EJR degree $c_{\emph{max}}$.
Proof.
We prove it by contradiction. Suppose the committee, $W$, output by $\lambda$-LS-PAV provides the EJR degree strictly smaller than $c_{\text{max}}$. There exists a $\ell$-cohesive group $V\subseteq N$, such that less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, i.e., |$|\{i\in V:|A_{i}\cap W|\geq\ell\}|<c_{\text{max}}$. Since $V$ is $\ell$-cohesive, there exist $\ell$ candidates approved by all voters in $V$. At least one such candidate, $c^{+}\in\bigcap_{i\in V}A_{i}$, is not selected, as otherwise all voters in $V$ approve at least $\ell$ members of $W$.
We will show that there exists a candidate $c^{-}\in W$ that the increment of the score by swapping $c^{+}$ and $c^{-}$ is at least $\lambda$, so $W$ cannot be an output of $\lambda$-LS-PAV (since the $while$ loop can be executed continually), a contradiction. To see this, we try to swap $c^{+}$ and any candidate $c^{-}\in W$. Since $c^{+}\in A_{i}$ for all $i\in V$, we have
$\Delta(W,c^{+},c^{-})$
$=$ $\sum_{\begin{subarray}{c}i:c^{+}\in A_{i}\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{\begin{subarray}{c}i:c^{+}\notin A_{i}\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
$\geq$ $\sum_{i\in V:c^{-}\notin A_{i}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V:c^{-}\in A_{i}}\frac{1}{|A_{i}\cap W|},$
and, by summing up $\Delta(W,c^{+},c^{-})$ for $c^{-}\in W$,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$
$\geq$ $\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in V:\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in N\setminus V:\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
$=$ $\sum_{i\in V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
$=$ $\sum_{i\in V}\frac{k-|A_{i}\cap W|}{|A_{i}\cap W|+1}-n+|V|=\sum_{i\in V}\frac{k+1}{|A_{i}\cap W|+1}-n$
$\geq$ $\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{|A_{i}\cap W|+1}-n\geq\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\!\frac{k+1}{\ell}-n.$
Since less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, the number of voters in $V$ approves less than $\ell$ members of $W$ is at least $|V|-c_{\text{max}}+1$. Thus,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$ $\geq\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}-n$
$\geq(k+1)\frac{|V|-c_{\text{max}}+1}{\ell}-n$
$\geq(k+1)(\frac{n}{k}-c_{\text{max}}+1)-n$
$=\frac{n}{k}-(k+1)(c_{\text{max}}-1)\geq\lambda k.$
From the pigeonhole principle it follows that there exists a candidate $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$. ∎
Since Proposition 3 holds for every initial committee $W$ in $\lambda$-LS-PAV. By considering the initial committee $W$ being the one with the maximum PAV-score and $\lambda=0$ (in which case the while-loop is never executed), we have the following corollary. In contrast to our counterexample in Appendix C, this corollary shows the success of PAV for large enough $n$, which may be of independent interest.
Corollary 1.
Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$. If $n\geq k(k+1)(c_{\emph{max}}-1)$, PAV has an EJR degree of $c_{\emph{max}}$.
By setting $\lambda=\frac{n}{k(k+1)}$ in Proposition 3, we have the following corollary, which is crucial for our algorithm.
Corollary 2.
Given any instance $(N,C,\mathbf{A},k)$ with maximum achievable EJR degree $c_{\emph{max}}$. If $n\geq k(k+1)^{2}(c_{\emph{max}}-1)$, $\frac{n}{k(k+1)}$-LS-PAV outputs a committee with EJR degree $c_{\emph{max}}$.
Based on Corollary 2, we can use a similar way to MDJR to design our algorithm. In particular, we can decide whether there exists a committee providing EJR degree $c$ by enumerating all the committees when $n\leq k(k+1)(c_{\text{max}}-1)$, which can be computed in running time $f(k,c)$. Our algorithm is presented in Algorithm 4, which achieves the maximum EJR degree and runs in time $f(k,c_{\text{max}})\cdot\text{poly}(m,n)$. The correctness of our algorithm follows from arguments similar to MDJR. The time complexity analysis is also similar, except that we need to show $\frac{n}{k(k+1)}$-LS-PAV runs in polynomial time. To see this, each swap operation can be executed in polynomial time with a minimum score increment by $\lambda=\frac{n}{k(k+1)}$, and the maximum PAV-score is $n\cdot(1+\frac{1}{2}+\cdots+\frac{1}{k})=O(n\ln k)$. Thus, the number of while-loop executions is bounded by $O(k^{2}\ln k)$.
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k)$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow\frac{n}{k(k+1)}\text{-LS-PAV}(\mathcal{I})$
2for $c:\lceil\frac{n}{k(k+1)}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
3       Enumerate all those $\binom{m}{k}$ possible committees to see if there is a committee $W^{*}$ achieving EJR degree $c$;
4       if $W^{*}$ exists then
5             $W\leftarrow W^{*}$
6      else
7             return $W$
8
Algorithm 4 MDEJR Voting Rule
4 Conclusion and Future Work
We initialize the study of the (E)JR degree and study its computational complexity and approximability. The (E)JR degree describes (E)JR from a quantitative perspective, which can help us better compare different committees. When explaining (E)JR from a stability perspective, i.e., for any $\ell$-cohesive group, if this group deviates and constructs $\ell$ winners, then at least one member does not want to deviate, as the current satisfaction is already $\ell$, which is the maximum satisfaction possible with $\ell$ winners. However, in reality, if only one person is represented, they can easily be persuaded to deviate by the rest of the group. Hence, if a committee provides a larger (E)JR degree, the possibility of deviation will be reduced. Moreover, we give complete pictures of both optimization problems, from a general NP-hardness with almost tight inapproximability to a parameterized complexity analysis with some natural parameters.
Many potential further works can be explored. For example, one can explore whether the negative results can be circumvented by considering restricted domains of preferences [13, 37]. Another direction is to consider some other quantitative measurements with respect to (E)JR, e.g., using the ratio instead of the number. As we have discussed earlier, our definition of EJR degree using numbers is more aligned with the original definition of EJR, whereas the definition using ratios/fractions, by prioritizing more on $\ell$-cohesive groups with large $\ell$, is more aligned with the notion of Individual Representation [7] discussed in Sect. 1.3. We believe both the “number version” and the “ratio version” are worth-studied. Which choice is better depends on the specific applications.
Appendix A Parameterized Complexity Basics
A parameterized problem is a language $L\subseteq\Sigma^{\ast}\times\mathbb{Z}^{+}$ (where $\Sigma$ is the alphabet set) and $p$ in an instance $(x,p)\in\Sigma^{\ast}\times\mathbb{Z}^{+}$ is called the parameter. A parameterized problem is fixed-parameter tractable (FPT), or a parameterized problem is in the complexity class FPT, if there exists an algorithm that decides the problem with running time bounded by $f(p)\cdot\text{poly}(|(x,p)|)$, where $(x,p)$ is the instance with length $|(x,p)|$, poly()($\cdot$)( ⋅ ) is a polynomial function, and $f(\cdot)$ is a computable function. The definition of parameterized problems and FPT can be straightforwardly generalized to optimization problems.
As introduced above, the instance of a parameterized problem consists of two parts: the “original instance” $x$ and the parameter $p$. In our case of MDJR, $x$ corresponds to a MDJR instance $(N,C,\mathbf{A},k)$, and $p$ is a parameter that we need to specify. For example, in Section 3.1, we select the winner committee size $k$ as the parameter. If there is an algorithm that solves MDJR problem with running time $f(k)\cdot\text{poly}(m,n)$, then we will say that MDJR is fixed-parameter tractable parameterized by $k$. (However, as we will see later, this is unlikely the case.) Here, $f(k)$ can be any computable functions such as $f(k)=2^{k}$ or $f(k)=2^{2^{k}\cdot k!}$. Specifically, the running time is not required to be polynomial in terms of the parameter, as long as the (possibly) super-polynomial factor with the parameter in the time complexity can be separated from the polynomial function of the length of the original instance. Naturally, for the algorithm to be useful in practice, we should choose parameters that are not large in practical scenarios, such as the size of the committee (which is usually significantly smaller than the number of voters).
It is also straightforward to generalize the parameterized problems with more than one parameter. For example, with two parameters $p_{1}$ and $p_{2}$, the problem is in FPT if there exists an algorithm with running time bounded by $f(p_{1},p_{2})$ times a polynomial function of the input length, where $f(\cdot,\cdot)$ now becomes a computable function of two variables. In Section 3.2, we consider the choice of two parameters: the committee size $k$ and the maximum achievable JR degree $c_{\text{max}}$, and we will show MDJR is FPT for this choice.
Downey and Fellows define the class of parameterized problems, the W-hierarchy, in an attempt to capture hard problems. Problems that are complete in the classes of W-hierarchy (such as W[1], W[2], etc) are believed to be fixed-parameter intractable. For the purpose of this paper, we omit the detailed definition of the W-hierarchy, and we refer the readers to Reference [11] for details.
To show a problem $L$ is W[$t$]-hard, we can give a parameterized reduction from a known W[$t$]-hard problem $\bar{L}$. We say that $\bar{L}$ parameterized reduce to $L$ if, given an $\bar{L}$ instance $(x,p)$, we can construct an $L$ instance $(x^{\prime},p^{\prime})$ such that
  1. 1.
    $(x,p)$ is a yes instance if and only if $(x^{\prime},p^{\prime})$ is a yes instance;
  2. 2.
    $p^{\prime}\leq g(p)$ for some computable function $g$;
  3. 3.
    the construction can be done in $f(k)\cdot\text{poly}(|x|)$ time for some computable function $f$.
It is straightforward to verify that, if $\bar{L}$ parameterized reduces to $L$, then $L$ being fixed-parameter tractable implies that $\bar{L}$ is fixed-parameter tractable.
The notion of W[$t$]-hardness can be extended to optimization problems, just as we can use the word “NP-hardness” to describe an optimization problem. Correspondingly, taking the maximization problem for example, description 1 above is changed such that a yes instance $(x,p)$ is mapped to instances $(x^{\prime},p^{\prime})$ with the objective value of at least a certain threshold $\theta$ while a no instance $(x,p)$ is mapped to instances with objective values less than $\theta$.
Appendix B Proof of Proposition 1
Fix $\gamma>0$. We construct the following ballot instance. Let $P$ be a large integer whose value is decided later, and let $L=2P^{2}-2P$. The instance consists of $n=P^{2}(L+2)-2P-L$ voters and $2(L+2)$ candidates $c_{1},\ldots,c_{2(L+2)}$. The winner committee has size $k=2L$. The ballot is described in the table below.
candidates voters who approve the candidates
$c_{1},c_{2}$ $1,2,\ldots,P^{2}$
$c_{3},c_{4}$ $P^{2},P^{2}+1,\ldots,2P^{2}-1$
$c_{5},c_{6}$ $22P^{2}-1,2P^{2},\ldots,3P^{2}-$ 2
$c_{7},c_{8}$ $33P^{2}-2,3P^{2}-1,\ldots,4P^{2}-$ 3
$\vdots$ $\vdots$
$c_{2L+1},c_{2L+2}$ $Lp^{2}-(L-1),\ldots,(L+1)P^{2}-L$
$c_{2L+3},c_{2L+4}$$\left\{\begin{array}[]{l}(L+1)P^{2}-L-P+1,\ldots,n\\ 1,2,\ldots,P\end{array}\right.$
We begin by identifying all cohesive groups. Let $V_{i}$ be the set of voters who approve the two candidates $c_{2i-1},c_{2i}$. We have $|V_{1}|=|V_{2}|=\cdots=|V_{L+2}|=P^{2}$: this is easy to see for $V_{1},\ldots,V_{L+1}$; for $V_{L+2}$, we have
$|V_{L+2}|=n-((L+1)P^{2}-L-P+1)+1+P,$
which equals to $P^{2}$ by substituting the value of $n=P^{2}(L+2)-2P-L$. In addition, we have $P^{2}=2\cdot\frac{n}{k}$ (to see this equation, substitute the values of $n=P^{2}(L+2)-2P-L$, $L=2P^{2}-2P$, and $k=2L$), so all of $V_{1},\ldots,V_{L+2}$ are $2$-cohesive groups, and it is easy to see that these are the only $2$-cohesive groups. Those $1$-cohesive groups are exactly those subsets of each $V_{i}$ with at least $\frac{n}{k}=\frac{P^{2}}{2}$ voters. We have identified all the cohesive groups.
By our construction, we have
$|V_{i}\cap V_{j}|=\left\{\begin{array}[]{ll}1&\mbox{if }j=i+1\mbox{ and }j\leq L+1\\ P&\mbox{if }i=L+1\mbox{ and }j=L+2\\ P&\mbox{if }i=L+2\mbox{ and }j=1\\ 0&\mbox{otherwise}\end{array}\right..$
To maximize the EJR degree, it is easy to see that the optimal winner committee can only be $\{c_{3},c_{4},\ldots,c_{2L},c_{2L+3},c_{2L+4}\}$, in which case only voters in $V_{1}$ and $V_{L+1}$ (respectively) are not fully represented. The optimal EJR degree is $P+1$.
To maximize the JR degree, we need to select at least one candidate from $\{c_{2i-1},c_{2i}\}$ for each $i=1,\ldots,L+2$, in which case all voters have at least $1$ approved candidate. The optimal JR degree is the maximum possible value $\frac{n}{k}=\frac{P^{2}}{2}$. The remaining $L-2$ candidates can be selected arbitrarily.
To see 1 in the proposition, we have $c^{\ast}_{\text{JR}}/c^{\ast}_{\text{EJR}}=\frac{P^{2}}{2P+2}$, which can be made larger than $\gamma$ by large enough $P$.
To see 2, the described winner committee optimal for the EJR degree has a JR degree of $P+1$, which is less than $c^{\ast}_{\text{JR}}/\gamma$ for large enough $P$.
To see 3, for a winner committee that optimizes the JR degree, among the $L+2$ candidate groups $\{c_{1},c_{2}\},\ldots,\{c_{2L+3},c_{2L+4}\}$, there are exactly four of them with only one selected candidate. By the pigeonhole principle, there is at least one of the $L-1$ candidate groups $\{c_{3},c_{4}\},\{c_{5},c_{6}\},\ldots,\{c_{2L-1},c_{2L}\}$, with one selected candidate. Since $V_{i}$ intersects with $\bigcup_{j\neq i}V_{j}$ by only two elements for each $i=2,\ldots,L$, the EJR degree in this case is at most $2$, which is less than $c^{\ast}_{\text{E1JR}}/\gamma$ for large enough $P$.
Appendix C Counterexample for PAV failing MDEJR
Consider an instance with $m=3p+2$ candidates $c_{1},c_{2},\dots,c_{3p},d_{1},d_{2}$ and $n=3p+(3p+2)+p+(9p^{2}-p-1)=(3p+1)^{2}$ voters. The committee size is set to $k=3p+1$. There are $4$ groups of voters, $D_{1},D_{2},T$, and $S$, with $|D_{1}|=3p$, $|D_{2}|=3p+2$, $|T|=p$, and $|S|=9p^{2}-p-1$. All voters in $D_{1}$ approve $d_{1}$ and $d_{2}$, and the $i$-th voter in $D_{1}$ approves $c_{i}$ additionally. All voters in $D_{2}$ approve $d_{1}$ and $d_{2}$. For those $p$ voters in group $T$, the $i$-th voter approves $c_{3i-2},c_{3i-1},c_{3i}$. All voters in $S$ approve $c_{1},c_{2},\dots,c_{3p}$.
Note that any size-$k$ committee covers every voter, so the JR-degree is always $\frac{n}{k}=3p+1$. We only need to care about $\ell$-cohesive groups with $\ell\geq 2$. First, any $\ell$-cohesive group with $\ell\geq 4$ is a subset of $S$ since voters in the other three group have degree less than $4$. There is no $\ell$-cohesive group for $\ell\geq k-2=3p-1$ as $(3p-1)(3p+1)=9p^{2}-1>|S|$.
At least $13p-$ 1 candidates in $c_{1},\dots,c_{3p}$ are selected, so any $\ell$-cohesive group with $\ell\geq 2$ has $\ell\frac{n}{k}>\frac{n}{k}3p+1$ voters that approve at least $\ell$ candidates in the committee.
Any voter in $T$ together with some voters in $S$ forms a $2$-cohesive group or $3$-cohesive group, has at least $13p+$ 1 voters in which approve at least $13p+$ 1 candidates in the committee.
In addition, $D_{1}\cup D_{2}$ forms a $2$-cohesive group. Since there is only one voter (in $D_{1}$) approves some $c_{i}$ and $d_{1}$ (or $d_{2}$), $D_{1}\cup D_{2}$ is the only $\ell$-cohesive group that contains voters in $D_{1}$ and $D_{2}$. Thus, the EJR-degree of the committee depends on the number of voters in $D_{1}\cup D_{2}$ that approve at least $2$ candidates in the committee. There are only two kinds of different committees:
Case 1
If we select any $13p-$ 1 candidates in $\{c_{1},\dots,c_{3p}\}$ and $d_{1},d_{2}$ to form a winning committee, all voters in $D_{1}\cup D_{2}$ approve at least $2$ winning candidates, namely, $d_{1}$ and $d_{2}$. Therefore, the EJR degree reaches its maximum, $13p+$ 1.
Case 2
If we select all $3p$ candidates in $\{c_{1},\dots,c_{3p}\}$ and one of $d_{1},d_{2}$, the EJR-degree is $3p$. To see this, for the $2$-cohesive group $D_{1}\cup D_{2}$, all voters in $D_{2}$ only approve $1$ candidate in the winner committee, so the number of voters who approve $2$ winner candidates in this $2$-cohesive group is $3p$.
Consider the winner committee $W=\{c_{1},\dots,c_{3p},d_{1}\}$ in Case $2$. If we swap $d_{2}$ and some $c_{i}$, the PAV-score increases by $\frac{1}{3}(|D_{1}|-1)$ due to voters in $D_{1}$, by $\frac{1}{2}|D_{2}|$ due to voters in $D_{2}$, while the PAV-score decreases by $\frac{1}{3}$ for voters in $T$ and by $\frac{1}{3p}|S|$ for voters in $S$. Therefore, the change in the PAV-score is
$\Delta(W,d_{2},c_{i})$ $=\frac{1}{3}(3p-1)+\frac{1}{2}(3p+2)-\frac{1}{3}-\frac{9p^{2}-p-1}{3p}$
$=-\left(\frac{1}{2}p-\frac{2}{3}-\frac{1}{3p}\right).$
For all $p\geq 2$, $\frac{1}{2}p-\frac{2}{3}-\frac{1}{3p}>0$, which implies Case $2$ has a larger PAV-score compared with Case $1$. The voting rule PAV outputs the committee in Case $2$, which fails to maximize the EJR degree.

### v2

@@
 Justified Representation (JR)/Extended Justified Representation (EJR) is a desirable axiom in multiwinner approval voting.
-In contrast to (E)JR only requires at least one voter to be represented in every cohesive group, we study its optimization version that maximizes the number of represented voters in each group.
-Given an instance, we say a winning committee provides an (E)JR degree of $c$ if at least $c$ voters in each $\ell$-cohesive group have approved $\ell$ winning candidates.
+In contrast to that (E)JR only requires at least one voter to be represented in every cohesive group, we study its optimization version that maximizes the number of represented voters in each group.
+Given an instance, we say a winning committee provides a JR degree (EJR degree, resp.) of $c$ if at least $c$ voters in each $\ell$-cohesive group ($1$-cohesive group, resp.) have approved $\ell$ ($1$, resp.) winning candidates.
 Hence, every (E)JR committee provides the (E)JR degree of at least $1$.
@@
 An (E)JR committee, which can be found in polynomial time, straightforwardly gives a $(k/n)$-approximation.
-On the other hand, we show that it is NP-hard to approximate MDJR and MDEJR to within a factor of $\left(k/n\right)^{1-\epsilon}$, for any $\epsilon>0$, which complements the approximation.
+We also show that the original algorithms for finding a JR and an EJR winner committee are also $1/k$ and $1/(k+1)$ approximation algorithms for MDJR and MDEJR respectively.
+On the other hand, we show that it is NP-hard to approximate MDJR and MDEJR to within a factor of $\left(k/n\right)^{1-\epsilon}$ and to within a factor of $(1/k)^{1-\varepsilon}$, for any $\epsilon>0$, which complements the positive results.
 Next, we study the fixed-parameter-tractability of this problem.
@@
 Definition 2 (Extended Justified Representation).
-[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies EJR for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there are at least one voter which approves at least $\ell$ candidates in $W$.
+[2] Given a ballot profile $\mathbf{A}=(A_{1},\dots,A_{n})$ over a candidate set $C$ and a committee size $k$, we say that a set of candidates $W$ of size $|W|=k$ satisfies EJR for $(\mathbf{A},k)$ if, for every $\ell$-cohesive group with every $\ell\in[k]$, there is at least one voter which approves at least $\ell$ candidates in $W$.
 We say that an ABC rule satisfies EJR if for each profile $\mathbf{A}$ and committee size $k$, each winning committee provide EJR.
-While there is a wealth of concepts associated with JR, there are instances where comparing the performance of different winning committees becomes impossible.
+While there are a wealth of concepts associated with JR, there are instances where comparing the performance of different winning committees becomes impossible.
 To illustrate this situation, consider a brief example.
@@
   • $c_{3}$ is approved by voters $7,8,9,1,2$;
-  • each of voters $c_{4},c_{5},c_{6}$ is approved by voters $1,2,4,5,7,8$.
+  • each of the candidates $c_{4},c_{5},c_{6}$ is approved by voters $1,2,4,5,7,8$.
 In this example, the winner committee that optimizes the proportionality degree is $\{c_{4},c_{5},c_{6}\}$, where we have $f(1)=2$ and $f(2)=3$, and the value of $f$ is maximized at both $1$ and $2$ (where $f$ is the function defined in Definition 3 but with a fixed instance).
@@
 In this paper, we focus on the computational complexity and the approximability of the optimization problems MDJR and MDEJR. Our results are listed below.
+  • We first show that the algorithm for finding a JR committee proposed by Aziz et al.
+[2] also provides a $\frac{1}{k}$-approximation to MDJR, and the algorithm for finding an EJR committee proposed by Aziz et al.
+[3] also provides a $\frac{1}{k+1}$-approximation to MDEJR.
+On the other hand, since the maximum (E)JR degree for both MDJR and MDEJR is $\lceil\frac{n}{k}\rceil$, the approximation guarantees for the two algorithms above can also be written as $1/\lceil\frac{n}{k}\rceil\approx k/n$.
+To complement these positive results, we show almost tight inapproximability results.
 We show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(k/n)^{1-\epsilon}$ for any $\epsilon>0$.
-This complements the result that an EJR committee (i.e., winner committee with both JR degree and EJR degree at least $1$) always exists and can be computed in polynomial time, which implies a $(k/n)$-approximation for both MDJR and MDEJR.
+We also show that it is NP-hard to approximate MDJR (MDEJR) within a factor of $(1/k)^{1-\epsilon}$ for any $\epsilon>0$.
   • We study the fixed-parameter-tractability of this problem.
@@
 2 Inapproximability of MDJR and MDEJR
-We have seen in the previous section that an EJR winner committee always exists and can be computed in polynomial time.
-Since an EJR winner committee achieves a degree of at least $1$ for both JR and EJR and the maximum possible degree is $n/k$, this implies a $(k/n)$-approximation algorithm for both MDJR and MDEJR problems.
-To complement this positive result, we will show that both MDJR and MDEJR are NP-hard to approximate to within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
+In this section, we study the approximability of MDJR and MDEJR. We provide almost tight approximability of both problems, in terms of both $k$ and $k/n$.
+As we mentioned before, the maximum degree for both MDJR and MDEJR is $\lceil\frac{n}{k}\rceil$.
+On the other hand, any winner committee satisfying JR (EJR, resp.) gives a JR degree (EJR degree, resp.) of at least $1$.
+Therefore, the algorithm for finding a JR (EJR, resp.) committee provides a $1/\lceil\frac{n}{k}\rceil\approx\frac{k}{n}$ approximation to MDJR (MDEJR, resp.).
+In Sect.
+2.1, we will show that the algorithm for finding a JR committee proposed by Aziz et al.
+[2] also provides a $\frac{1}{k}$-approximation to MDJR, and the algorithm for finding an EJR committee proposed by Aziz et al.
+[3] also provides a $\frac{1}{k+1}$-approximation to MDEJR.
+In Sect. 2.2, we show that the approximation ratios of approximately $k/n$ and $1/k$ mentioned above are almost tight.
+2.1 Approximation Algorithms
+The algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2] always outputs a JR committee. We will show in Theorem 1 that it provides a $\frac{1}{k}$-approximation to MDJR.
+Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
+Output: A winning committee $W$ of size $k$
+1 $W\leftarrow\emptyset$
+2for $j\in[k]$ do
+3       Let $c$ be the candidate approved by the maximum number of voters in $N$
+4      $W\leftarrow W\cup c$
+5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
+return $W$
+Algorithm 1 Greedy Approval Voting (GreedyAV)
+We prove the following proposition first.
+Proposition 2.
+GreedyAV outputs a committee achieving JR degree at least $\frac{n}{k^{2}}$.
+Proof.
+We prove it by contradiction.
+Suppose the committee output by GreedyAV provides the JR degree less than $\frac{n}{k^{2}}$.
+Hence, there will be at least $(\frac{n}{k}-\frac{n}{k^{2}}+1)$ voters approving no candidate in $W$ that approve a common candidate $c\notin W$.
+From the definition of GreedyAV, the coverage of voters is at least $(\frac{n}{k}-\frac{n}{k^{2}}+1)$ in each iteration, and at least $\frac{n}{k}$ in the first iteration (given that there is at least one cohesive group).
+Therefore, the total number of voters is at least
+$\frac{n}{k}+(k-1)\left(\frac{n}{k}-\frac{n}{k^{2}}+1\right)+\left(\frac{n}{k}-\frac{n}{k^{2}}+1\right)=n+\frac{n}{k}-k\left(\frac{n}{k^{2}}-1\right)>n,$
+which leads to a contradiction in the number of voters. ∎
+Now, we are ready to show the approximation guarantee of GreedyAV.
+Theorem 1.
+GreedyAV runs in polynomial time and provides a $\frac{1}{k}$-approximation to MDJR.
+Proof.
+It is clear that the algorithm runs in polynomial time.
+To show the approximation guarantee, we show that the minimum JR degree of $\frac{n}{k^{2}}$ proved in Proposition 2.
+To see this, notice that we have remarked that the maximum possible JR degree is $\lceil\frac{n}{k}\rceil$.
+On the other hand, since the JR degree is an integer, the minimum JR degree that Greedy AV can guarantee is actually $\lceil\frac{n}{k^{2}}\rceil$.
+It suffices show that
+$\left\lceil\frac{n}{k^{2}}\right\rceil\geq\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil.$
+The inequality clearly holds if $n$ is a multiple of $k$.
+Suppose otherwise.
+Let $n=ak+b$ for non-negative integers $a,b$ with $11\leq b\leq k-$ 1.
+Let $a=sk+t$ for non-negative integers $s,t$ with $10\leq t\leq k-$ 1.
+For the left-hand side of the inequality, we have
+$\left\lceil\frac{n}{k^{2}}\right\rceil=\left\lceil\frac{(sk+t)k+b}{k^{2}}\right\rceil\geq\left\lceil s+\frac{b}{k^{2}}\right\rceil\geq s+1,$
+where the first inequality above is due to $t\geq 0$ and the second inequality above is due to $b\geq 1$. For the right-hand side of the inequality, we have
+$\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil=\frac{1}{k}\cdot(a+1)=\frac{sk+t+1}{k}\leq s+1,$
+where the last inequality is due to that $t\leq k-1$.
+Putting together, we have $\left\lceil\frac{n}{k^{2}}\right\rceil\geq\frac{1}{k}\cdot\left\lceil\frac{n}{k}\right\rceil$, which implies the theorem. ∎
+Similar to MDJR, we first consider an EJR voting rule, proportional approval voting (PAV) [2], which outputs the committee that maximizes the PAV-score, where the PAV-score of a committee $W\subseteq C$ is defined as
+$s_{\text{PAV}}(W)=\sum_{i=1}^{n}\sum_{j=1}^{|A_{i}\cap W|}\frac{1}{j}.$
+In the PAV-score, each voter’s “utility” is defined by the harmonic progression $H[t]$ for $t$ approved winning candidates and the PAV-score can then be understood as the social welfare.
+One may wonder whether PAV can provide the maximum EJR degree.
+We find that PAV fails to achieve the maximum EJR degree in some instances.
+A counterexample can be found in Appendix C.
+In addition, PAV cannot be computed in polynomial time.
+Aziz et al.
+[3] showed that a local search alternative algorithm for PAV can both satisfy EJR and be computed in polynomial time.
+The algorithm is described in Algorithm 2.
+Starting from an arbitrary winning committee, the algorithm considers all possible single-candidate-replacements that increase the PAV score by at least $\lambda$ (where $\lambda$ is a parameter of the algorithm).
+For each pair of candidate $c^{+}$ and $c^{-}$ with $c^{+}\notin W$ and $c^{-}\in W$, if we swap $c^{+}$ and $c^{-}$, i.e.
+to remove $c^{-}$ from the committee and select $c^{+}$ instead, the score is increased by $\Delta(W,c^{+},c^{-})=s_{\text{PAV}}(W\setminus\{c^{-}\}\cup\{c^{+}\})-s_{\text{PAV}}(W).$
+Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
+Output: A winning committee $W$ of size $k$
+1 $W\leftarrow k$ arbitrary candidates from $C$
+2while there exist $c^{+}\notin W$ and $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$ do
+3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
+return $W$
+Algorithm 2 $\lambda$-LS-PAV
+Next, we will show that $\lambda$-LS-PAV runs in polynomial time and provides a $\frac{1}{k+1}$-approximation to MDEJR for a suitable choice of $\lambda$.
+We show the following proposition first.
+Proposition 3.
+$\lambda$-LS-PAV outputs a committee achieving EJR degree at least $c^{\ast}=\frac{n}{k(k+1)}-\lambda\frac{k}{k+1}$ for any $\lambda\in[0,\frac{n}{k^{2}})$.
+Proof.
+We prove it by contradiction.
+Suppose the committee, $W$, output by $\lambda$-LS-PAV provides the EJR degree strictly smaller than $c^{\ast}$.
+There exists a $\ell$-cohesive group $V\subseteq N$, such that less than $c^{\ast}$ voters in $V$ approves at least $\ell$ members of $W$, i.e., |$|\{i\in V:|A_{i}\cap W|\geq\ell\}|<c^{\ast}$.
+Since $V$ is $\ell$-cohesive, there exist $\ell$ candidates approved by all voters in $V$.
+At least one such candidate, $c^{+}\in\bigcap_{i\in V}A_{i}$, is not selected, as otherwise all voters in $V$ approve at least $\ell$ members of $W$.
+We will show that there exists a candidate $c^{-}\in W$ such that the increment of the score by swapping $c^{+}$ and $c^{-}$ is at least $\lambda$, so $W$ cannot be an output of $\lambda$-LS-PAV, leading to a contradiction.
+To see this, we try to swap $c^{+}$ and any candidate $c^{-}\in W$.
+Since $c^{+}\in A_{i}$ for all $i\in V$, we have
+$\Delta(W,c^{+},c^{-})=$ $\sum_{\begin{subarray}{c}i:c^{+}\in A_{i}\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{\begin{subarray}{c}i:c^{+}\notin A_{i}\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
+$\geq$ $\sum_{i\in V:c^{-}\notin A_{i}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V:c^{-}\in A_{i}}\frac{1}{|A_{i}\cap W|},$
+and, by summing up $\Delta(W,c^{+},c^{-})$ for $c^{-}\in W$,
+$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})\geq$ $\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in V:\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in N\setminus V:\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
+$=$ $\sum_{i\in V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
+$=$ $\left(\sum_{i\in V}\frac{k-|A_{i}\cap W|}{|A_{i}\cap W|+1}\right)-\left(n-|V|\right)$
+$=$ $\left(\sum_{i\in V}\frac{k+1}{|A_{i}\cap W|+1}\right)-n$
+$\geq$ $\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{|A_{i}\cap W|+1}\right)-n$
+$\geq$ $\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}\right)-n.$
+Since less than $c^{\ast}$ voters in $V$ approve at least $\ell$ members of $W$, the number of voters in $V$ approve less than $\ell$ members of $W$ is at least $|V|-c^{\ast}+1$. Thus,
+$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$ $\geq\left(\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}\right)-n$
+$\geq(k+1)\frac{|V|-c^{\ast}+1}{\ell}-n$
+$>(k+1)\left(\frac{|V|}{\ell}-c^{\ast}\right)-n$
+$\geq(k+1)\left(\frac{n}{k}-c^{\ast}\right)-n$
+$=\frac{n}{k}-(k+1)c^{\ast}$
+$=\lambda k.$
+From the pigeonhole principle, it follows that there exists a candidate $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$. ∎
+Now, we are ready to conclude the approximation guarantee of $\lambda$-LS-PAV.
+Theorem 2.
+For $\lambda=\frac{1}{2k^{2}}$, $\lambda$-LS-PAV runs in polynomial time and provides a $\frac{1}{k+1}$-approximation to MDEJR.
+Proof.
+Firstly, the algorithm runs in polynomial time: the maximum possible PAV-score is $n\cdot\sum_{j=1}^{k}\frac{1}{j}=O(n\log k)$.
+On the other hand, the PAV-score is increased by at least $\frac{1}{2k^{2}}$.
+The algorithm terminates in at most $O(nk^{2}\log k)$ iterations, which runs in polynomial time.
+To prove the approximation guarantee, Proposition 3 shows that the EJR degree of the output committee is at least $\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}$.
+Since the degree is an integer, this is at least $\lceil\frac{n}{k(k+1)}\rceil$: if $\frac{n}{k(k+1)}$ is an integer, we have $\lceil\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}\rceil=\frac{n}{k(k+1)}$ as $\frac{1}{2k(k+1)}<1$; otherwise, since $\frac{n}{k(k+1)}-\lfloor\frac{n}{k(k+1)}\rfloor\geq\frac{1}{k(k+1)}>\frac{1}{2k(k+1)}$, the numbers $\frac{n}{k(k+1)}$ and $\frac{n}{k(k+1)}-\frac{1}{2k(k+1)}$ have the same integral part.
+Following a similar analysis as in the proof of Theorem 1, we can prove that
+$\left\lceil\frac{n}{k(k+1)}\right\rceil\geq\frac{1}{k+1}\cdot\left\lceil\frac{n}{k}\right\rceil,$
+which implies the theorem.
+Specifically, the inequality clearly holds if $k$ divides $n$.
+Otherwise, let $n=ak+b$ for integers $a\geq 0$ and $11\leq b\leq k-$ 1.
+Let $a=s(k+1)+t$ for integers $s\geq 0$ and $0\leq t\leq k$.
+We have
+$\left\lceil\frac{n}{k(k+1)}\right\rceil=\left\lceil\frac{sk(k+1)+tk+b}{k(k+1)}\right\rceil\geq s+1\geq\frac{s(k+1)+t+1}{k+1}=\frac{1}{k+1}\cdot(a+1)=\frac{1}{k+1}\cdot\left\lceil\frac{n}{k}\right\rceil.\qed$
+2.2 Matching Inapproximability Results
+To complement the positive results mentioned in the previous section, we present the following inapproximability results.
 Theorem 3.
@@
 It is NP-hard to approximate MDEJR within a factor of $\left(\frac{k}{n}\right)^{1-\epsilon}$ for any $\epsilon>0$.
-We will simultaneously prove the two theorems by constructing a hard ballot instance that is used for proving both theorems.
+Theorem 5.
+It is NP-hard to approximate MDJR within a factor of $\left(\frac{1}{k}\right)^{1-\epsilon}$ for any $\epsilon>0$.
+Theorem 6.
+It is NP-hard to approximate MDEJR within a factor of $\left(\frac{1}{k}\right)^{1-\epsilon}$ for any $\epsilon>0$.
+We will simultaneously prove these four theorems by constructing a hard ballot instance that is used for all of them.
 We will make sure the instance we constructed has no $\ell$-cohesive group with $\ell>1$.
 Notice that the JR degree and the EJR degree for any committee are always the same for instances with only $1$-cohesive groups.
-2.1 Proof of Theorem 1 and 2
-Before we prove the theorem, we first introduce a NP-hard problem: sparse-SAT problem. One can find variations of SAT problem and a similar argument in [36].
+In addition, we have $n\approx k^{2}$ in our construction, so the factor $k/n$ is approximately $1/k$.
+Before we prove the theorems, we first introduce a NP-hard problem: sparse-SAT problem. One can find variations of SAT problem and a similar argument in Tovey [36].
 Definition 7 (sparse-SAT).
-Given a CNF formula $\phi$ that, for any variables $x$ and $y$, at most one clause contains both $x$ (or $\neg x$) and $y$ (or $\neg y$), decide if there is a value assignment to the variables to make $\phi$ true.
-To see the NP-hardness, it can be reduced from the SAT problem.
+Given a CNF formula $\phi$ that, for any pair of variables $x$ and $y$, at most one clause contains both $x$ (or $\neg x$) and $y$ (or $\neg y$), decide if there is a value assignment to the variables to make $\phi$ true.
+To see its NP-hardness, it can be reduced from the SAT problem.
 Start with any SAT instance.
 Without loss of generality, suppose each variable, $x$ or $\neg x$, appears in each clause at most once.
-For each variable $x$ that there are more than two clauses containing $x$ or $\neg x$ perform the following procedure: Suppose $x$ appears in $k$ clauses.
-Create $k$ new variables $x_{1},\dots,x_{k}$ and replace the $i$th occurrence of $x$ with $x_{i}$ (and $\neg x$ is replaced by $\neg x_{i}$, respectively), $i=1,\dots,k$.
-Append the clause $(x_{i}\vee\neg x_{i+1})$ for $i=1,\dots,k-1$ and the clause $(x_{k}\vee\neg x_{1})$.
-Note that, in the new instance, variable $x_{i}$ and $y_{j}$ appear in a clause only when the $i$th occurrence of $x$ and the $j$th occurrence of $y$ in $\phi$ are in the same clause, so the new instance satisfies the limitation of sparse-SAT problem.
+For each variable $x$ such that $x$ or $\neg x$ is contained in more than two clauses, we perform the following procedure.
+Suppose $x$ appears in $k$ clauses.
+Create $k$ new variables $x_{1},\dots,x_{k}$ and replace the $i$th occurrence of $x$ with $x_{i}$ (and $\neg x$ is replaced by $\neg x_{i}$, respectively) for each $i=1,\dots,k$.
+Add the clause $(x_{i}\vee\neg x_{i+1})$ for $i=1,\dots,k-1$ and the clause $(x_{k}\vee\neg x_{1})$.
+Note that, in the new instance, variable $x_{i}$ and $y_{j}$ appear in a clause only when the $i$th occurrence of $x$ and the $j$th occurrence of $y$ in $\phi$ are in the same clause, so the new instance satisfies the requirement of the sparse-SAT problem.
 In the new instance, the clause $(x_{i}\vee\neg x_{i+1})$ implies that if $x_{i}$ is false, $x_{i+1}$ must be false as well.
@@
 Now we are ready to prove our theorems.
-Proof of Theorem 1 and 2.
+[…]
+Figure 1: An example of the construction with $\phi=(x_{1}\vee\neg x_{3})\wedge(\neg x_{1}\vee x_{2})$.
+Edges in the graph represent approvals.
+An edge connecting between a group of voters and a group of candidates indicates that every voter in the voter group approves every candidate in the candidate group.
+Proof of Theorem 3 and 4.
 We reduce from sparse-SAT problem.
-Given any sparse-SAT instance $\phi$, suppose there are $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$).
-We consider a ABC voting instance with $13\bar{m}+\bar{n}+$ 1 candidates, $c_{1},\dots,c_{2\bar{m}}$, $s_{1},\dots,s_{\bar{n}}$, $t_{1},\dots,t_{\bar{m}}$, $d$ and $\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}$ voters.
+Given any sparse-SAT instance $\phi$, suppose there are $\bar{n}$ clauses and $\bar{m}$ variables (say $x_{1},\dots,x_{\bar{m}}$).
+We consider an ABC voting instance with $13\bar{m}+\bar{n}+$ 1 candidates $\{c_{1},\dots,c_{2\bar{m}}$, $s_{1},\dots,s_{\bar{n}}$, $t_{1},\dots,t_{\bar{m}},d\}$ and $\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}$ voters.
 We want to select a committee of size $\bar{m}+1$.
 Hence, we care about the cohesive group of size $\frac{\bar{n}\bar{m}+\bar{m}^{2}+\bar{n}+\bar{m}}{\bar{m}+1}=\bar{n}+\bar{m}$.
-First, for each variable $x_{j}$ and its complement $\neg x_{j}$, we create two corresponding candidates $c_{2j-1}$ and $c_{2j}$ and a group $T_{j}$ of $\bar{m}$ voters who approve $c_{2j-1}$ and $c_{2j}$.
+First, for each variable $x_{j}$ and its negation $\neg x_{j}$, we create two corresponding candidates $c_{2j-1}$ and $c_{2j}$ and a group $T_{j}$ of $\bar{m}$ voters who approve $c_{2j-1}$ and $c_{2j}$.
 For the $i$th clause, we create a group $S_{i}$ of $\bar{m}$ voters.
@@
 We create a set $D$ of $\bar{n}$ voters who approve $s_{1},\dots,s_{\bar{n}},t_{1},\dots,t_{\bar{m}},d$.
-Hence, for each $i\in[\bar{n}]$, $S_{i}\cup D$ forms a $1$-cohesive group and for each $j\in[\bar{m}]$, $T_{j}\cup D$ forms a $1$-cohesive group.
+Hence, for each $i\in[\bar{n}]$, $S_{i}\cup D$ forms a $1$-cohesive group, and for each $j\in[\bar{m}]$, $T_{j}\cup D$ forms a $1$-cohesive group.
 Moreover, we create a set $D^{+}$ of $\bar{m}$ voters who approve $d$.
 Hence, $D\cup D^{+}$ forms a $1$-cohesive group.
+An example of our construction is shown in Fig.
+1.
 Notably, there is no $2$-cohesive group.
@@
 For candidates $c_{1},\dots,c_{2\bar{m}}$, we will show that no $2(\bar{n}+\bar{m})$ voters have two common approved candidates.
-Since in a sparse-SAT instance, for any two variables $x_{i}$ and $x_{j}$, at most one clause contains them or their complement, for any two candidates $c_{2i}$ or $c_{2i-1}$ and $c_{2j}$ or $c_{2j-1}$, at most one group $S_{k}$ of $\bar{m}$ voters approves them simultaneously.
-Moreover, only $\bar{m}$ voters in group $T_{i}$ approve both $c_{2i-1}$ and $c_{2i}$.
-So there is no $2$-cohesive group or $\ell$-cohesive group for $l>2$.
-In this case, the EJR degree equals to JR degree.
-If there is a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of $\bar{n}+\bar{m}$.
-To see this, we elect $c_{2j-1}$ as a winner if $x_{j}$ assigned true or we elect $c_{2j}$ for all $j\in[\bar{m}]$.
-Then we elect $d$ as the winner.
-We can verify that all voters in $S_{i}$ and $T_{j}$ are covered for every $i\in[\bar{n}],j\in[\bar{m}]$.
-In addition, $d$ covers all voters in $D$ and $D^{+}$.
-Hence, all voters are covered and the JR degree equals the size of $1$-cohesive group, $\bar{n}+\bar{m}$, and so does the EJR degree.
-If there does not exist a value assignment to the variables to make $\phi$ true, then MDEJR will achieve the EJR degree of at most $\bar{n}$.
+For any two candidates $t_{2i-1},t_{2i}$ in $\{c_{1},\dots,c_{2\bar{m}}\}$ that correspond to a variable and its negation, the set of voters approving both candidates is exactly $T_{i}$ (notice that we can assume without loss of generality that $x_{i}$ and $\neg x_{i}$ do not appear in the same clause, for otherwise, we can safely remove both literals).
+We have $|T_{i}|=\bar{m}<2(\bar{n}+\bar{m})$.
+For any two candidates in $\{c_{1},\dots,c_{2\bar{m}}\}$ that correspond to different variables, by our sparsity assumption of $\phi$, the set of voters approving both candidates is at most some $S_{j}$ (representing a clause that contains both variables).
+We have $|S_{j}|=\bar{m}<2(\bar{n}+\bar{m})$.
+Thus, there is no $2$-cohesive group or $\ell$-cohesive group for $\ell>1$.
+In this case, the EJR degree is equal to the JR degree.
+Thereafter, we will analyze the JR degree and MDJR, and the analysis applies to MDEJR as well.
+If there is a value assignment to the variables to make $\phi$ true, then MDJR will achieve a JR degree of $\bar{n}+\bar{m}$.
+To see this, for each $j\in[\bar{m}]$, we select $c_{2j-1}$ as a winner if $x_{j}$ is assigned true, or we select $c_{2j}$.
+Then we select $d$ as the winner.
+We can verify that each voter in $\{S_{i}\}_{i\in[\bar{n}]}\cup\{T_{j}\}_{j\in[\bar{m}]}$ approves at least one winner.
+In addition, $d$ is approved by all voters in $D$ and $D^{+}$.
+Hence, all voters approve at least one winner, and the JR degree equals the size of a $1$-cohesive group, which is $\bar{n}+\bar{m}$.
+If there does not exist a satisfying assignment to $\phi$, then MDJR will achieve the JR degree of at most $\bar{n}$.
 To see this, we prove it by contradiction.
-Assume that MDEJR can achieve the EJR degree larger than $\bar{n}$.
-If $d$ is not elected, no voter in $D^{+}$ can be covered since they only approve $d$.
-Hence, for cohesive group $D\cup D^{+}$, there are at most $\bar{n}$ voters can be covered, which makes a contradiction.
-If $d$ is elected, there will be $\bar{m}$ vacant seats.
-If voters in $T_{j}$ are not covered for $j\in[\bar{m}]$, there are at most $\bar{n}$ voters covered in cohesive group $D\cup T_{j}$, a contradiction, so all voters in group $T_{j}$ must be covered for all $j\in[\bar{m}]$, indicating that exact one of $3$ candidates $\{c_{2j-1},c_{2j},t_{j}\}$ is elected.
-Next, we will show that there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered.
-To see this, suppose there is a committee $W^{\prime}$ of size $\bar{m}$ which covers all groups $S_{1},\dots,S_{\bar{n}}$.
-For every $t_{j}\in W^{\prime}$, we can find $c_{2j-1}$ or $c_{2j}$ to replace $t_{j}$ since $t_{j}$ only covers group $T_{j}$.
-Hence, we can find one candidate in $\{c_{2i-1},c_{2i}\}$ for all $i\in[\bar{m}]$ to cover all groups, implying that there exists a value assignment to the variables ($x_{i}$ is assigned true if $c_{2i-1}$ is elected or false otherwise) to make $\phi$ true, leading to a contradiction.
-Therefore, there is some group among $S_{1},\dots,S_{\bar{n}}$ that cannot be covered.
+Assume that MDJR can achieve the JR degree of larger than $\bar{n}$ by the winner committee $W$ with $|W|=k=\bar{m}+1$.
+If $d\notin W$, no voter in $D^{+}$ can be covered since they only approve $d$.
+Hence, for the cohesive group $D\cup D^{+}$, at most $\bar{n}$ voters can be covered, which leads to a contradiction.
+Thus, $d\in W$.
+Let $W^{\prime}=W\setminus\{d\}$, and we have $|W^{\prime}|=\bar{m}$.
+If voters in $T_{j}$ are not covered for some $j\in[\bar{m}]$, at most $\bar{n}$ voters can be covered in the cohesive group $D\cup T_{j}$, leading to a contradiction.
+Thus, all voters in group $T_{j}$ must be covered for each $j\in[\bar{m}]$, indicating that exactly one of the $3$ candidates $\{c_{2j-1},c_{2j},t_{j}\}$ for each $j\in[\bar{m}]$ is selected (at least one candidate in each three-candidates group must be selected, we have $\bar{m}$ groups, and we have $|W^{\prime}|=\bar{m}$).
+Next, we will show that voters in at least one group among $S_{1},\dots,S_{\bar{n}}$ cannot be covered.
+Suppose this is not the case.
+For every $t_{j}\in W^{\prime}$, we can use either $c_{2j-1}$ or $c_{2j}$ to replace $t_{j}$ since $t_{j}$ only covers group $T_{j}$ and each of $c_{2j-1}$ and $c_{2j}$ covers at least $T_{j}$.
+Hence, we can find one candidate in $\{c_{2i-1},c_{2i}\}$ for all $i\in[\bar{m}]$ to cover all groups, implying that there exists a value assignment to the variables ($x_{i}$ is assigned true if $c_{2i-1}$ is selected, or false otherwise) to make $\phi$ true, leading to a contradiction.
+Therefore, at least one group among $S_{1},\dots,S_{\bar{n}}$ cannot be covered.
 Without loss of generality, we assume that it is $S_{i}$.
-Then, for cohesive group $S_{i}\cup D$, there are at most $\bar{n}$ voters that can be covered, which makes a contradiction.
-From the proof of NP-hardness, we further have MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}}{\bar{n}+\bar{m}}$ (Note that $\bar{n}$ and $\bar{m}$ are the parameters of the sparse-SAT problem rather than ABC voting, we will reformulate it later).
+Then, for the cohesive group $S_{i}\cup D$, at most $\bar{n}$ voters can be covered, leading to a contradiction.
+From the proof of NP-hardness, we have that MDJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}}{\bar{n}+\bar{m}}$, where $\bar{n}$ is the number of the clauses and $\bar{m}$ is the number of the variables.
 Now, instead of reducing from sparse-SAT problems directly, we add an intermediate reduction.
-Given a sparse-SAT instance $\phi$ with $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$), there is a corresponding ‘overstaffing’ sparse-SAT instance $\phi^{\prime}$ with $\bar{n}^{\prime}=\bar{n}+1$ clauses and $\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ different variables such that a new clause is added with a number of $(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ new variables like $(y_{1}\vee y_{2}\vee\dots)$.
-Obviously, the ‘overstaffing’ sparse-SAT problem is also an NP-hard problem.
-Now we reduce our problem from ‘overstaffing’ sparse-SAT problem in the same way that we reduce from sparse-SAT problem.
-Similarly, MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}$ where $\bar{n}^{\prime}=\bar{n}+1,\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$.
-Recall the construction, the number of voters and the committee size are $n_{\text{voting}}=\bar{n}^{\prime}\bar{m}^{\prime}+\bar{m}^{\prime 2}+\bar{n}^{\prime}+\bar{m}^{\prime}$ and $k_{\text{voting}}=\bar{m}^{\prime}+1$, respectively, which can be reformulated as $\frac{k_{\text{voting}}}{n_{\text{voting}}}=\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}$.
+Given a sparse-SAT instance $\phi$ with $\bar{n}$ clauses and $\bar{m}$ different variables, we construct another sparse-SAT instance $\phi^{\prime}$ with $\bar{n}^{\prime}=\bar{n}+1$ clauses and $\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ different variables such that a new clause is added with $(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ new variables.
+Obviously, $\phi$ and $\phi^{\prime}$ have the same satisfiability.
+Now, we use $\phi^{\prime}$ instead of $\phi$ to construct the ABC voting instance.
+By our previous analysis, MDJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}$ where $\bar{n}^{\prime}=\bar{n}+1,\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$.
+Recall that the number of voters and the committee size are $n_{\text{voting}}=\bar{n}^{\prime}\bar{m}^{\prime}+\bar{m}^{\prime 2}+\bar{n}^{\prime}+\bar{m}^{\prime}$ and $k_{\text{voting}}=\bar{m}^{\prime}+1$, respectively, which can be reformulated as $\frac{k_{\text{voting}}}{n_{\text{voting}}}=\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}$.
 Since
-$(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}=(\bar{n}^{\prime}+\bar{m}+(\bar{n}^{\prime}+\bar{m})^{\lceil\frac{1}{\epsilon}\rceil})^{\epsilon}>\bar{n}^{\prime}+\bar{m}>\bar{n}^{\prime}.$
-Finally, MDJR cannot be approximated in polynomial time to within a factor of
+$(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}=\left(\bar{n}^{\prime}+\bar{m}+(\bar{n}^{\prime}+\bar{m})^{\lceil\frac{1}{\epsilon}\rceil}\right)^{\epsilon}>\bar{n}^{\prime}+\bar{m}>\bar{n}^{\prime},$
+MDJR cannot be approximated in polynomial time to within a factor of
 $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}\leq\frac{(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}}{\bar{n}^{\prime}+\bar{m}^{\prime}}=\left(\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}\right)^{1-\epsilon}\!\!\!=\!\left(\frac{k_{\text{voting}}}{n_{\text{voting}}}\right)^{1-\epsilon},$
-where $\epsilon>0$. ∎
+where $\epsilon>0$.
+As we have mentioned, the same analysis holds for MDEJR, as there is no $\ell$-cohesive group with $\ell\geq 2$ in our construction. ∎
+Proof of Theorem 5 and 6.
+Both theorems can be proved with only minor modifications to the last step in the proof of Theorem 3 and 4 above:
+$\left(\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}\right)^{1-\epsilon}<\left(\frac{1}{1+\bar{m}^{\prime}}\right)^{1-\epsilon}=\left(\frac{1}{k_{\text{voting}}}\right)^{1-\epsilon}.\qed$
 3 Parameterized Complexity of MDJR/MDEJR
@@
 3.2.1 Algorithm for MDJR
-Our starting point is the algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2], which always outputs a JR committee.
-Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
-Output: A winning committee $W$ of size $k$
-1 $W\leftarrow\emptyset$
-2for $j\in[k]$ do
-3       Let $c$ be the candidate approved by the maximum number of voters in $N$
-4      $W\leftarrow W\cup c$
-5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
-return $W$
-Algorithm 1 Greedy Approval Voting (GreedyAV)
-Next, we show the following property for GreedyAV which is the key for our algorithm. It states that the algorithm GreedyAV also gives us the optimal JR degree if $n$ is large enough.
+Our starting point is the algorithm GreedyAV (Algorithm 1) proposed by Aziz et al.
+[2].
+We show the following property for GreedyAV which is the key for our algorithm.
+It states that the algorithm GreedyAV also gives us the optimal JR degree if $n$ is large enough.
 Proposition 4.
-Given any instance with the maximum achievable JR degree of $c_{\emph{max}}$ and $n>k^{2}(c_{\emph{max}}-1)$, GreedyAV will output the committee achieving JR degree $c_{\emph{max}}$.
+Given any instance with the maximum achievable JR degree of $c_{\max}$ and $n>k^{2}(c_{\max}-1)$, GreedyAV will output a committee achieving JR degree $c_{\max}$.
 Proof.
-We prove it by contradiction.
-Suppose the committee output by GreedyAV provides the JR degree less than $c_{\text{max}}$.
-Hence, there will be at least $(\frac{n}{k}-c_{\text{max}}+1)$ voters that approve a non-elected candidate.
-From the definition of GreedyAV, the coverage of voters is at least $(\frac{n}{k}-c_{\text{max}}+1)$ in each iteration, and at least $\frac{n}{k}$ in the first iteration.
-Therefore, the total number of voters is at least
-$\frac{n}{k}+(k-1)\left(\frac{n}{k}-c_{\text{max}}+1\right)+\left(\frac{n}{k}-c_{\text{max}}+1\right)$
-$=n+\frac{n}{k}-k(c_{\text{max}}-1)>n,$
-which leads to a contradiction in the number of voters. ∎
+$n>k^{2}(c_{\max}-1)$ implies $c_{\max}<\frac{n}{k^{2}}+1$. This proposition then follows from Proposition 2. ∎
 Given any instance with $n>k^{2}(c_{\text{max}}-1)$, GreedyAV can achieve the maximum JR degree.
 Hence, the remaining case is $n\leq k^{2}(c_{\text{max}}-1)$.
-Given any instance $(N,C,\mathbf{A},k)$, each candidate $c_{i}\in C$ can be seen as a subset of $N$, including all voters that approve $c_{i}$.
+Given any instance $(N,C,\mathbf{A},k)$, each candidate $c_{i}\in C$ can be seen as a subset of $N$ including all voters that approve $c_{i}$.
 Hence, there are at most $2^{n}$ different types of candidates, implying that every instance corresponds to an equivalent instance with $m\leq 2^{n}$.
@@
 3.2.2 Algorithm for MDEJR
-Similar to MDJR, we first consider an EJR voting rule, proportional approval voting (PAV) [2], which outputs the committee that maximizes the PAV-score, where the PAV-score of a committee $W\subseteq C$ is defined as
-$s_{\text{PAV}}(W)=\sum_{i=1}^{n}\sum_{j=1}^{|A_{i}\cap W|}\frac{1}{j}.$
-In the PAV-score, each voter’s “utility” is defined by the harmonic progression $H[t]$ for $t$ approved winning candidates and the PAV-score can then be understood as the social welfare.
-One may wonder whether PAV can provide the maximum EJR degree.
-We find that PAV fails to achieve the maximum EJR degree in some instances.
-A counterexample can be found in Appendix C.
-In addition, PAV cannot be computed in polynomial time.
-Aziz et al.
-[3] showed that a local search alternative algorithm for PAV can both satisfy EJR and be computed in polynomial time.
-The algorithm is described in Algorithm 3.
-Starting from an arbitrary winning committee, the algorithm considers all possible single-candidate-replacements that increase the PAV score by at least $\lambda$ (where $\lambda$ is a parameter of the algorithm).
-For each pair of candidate $c^{+}$ and $c^{-}$ with $c^{+}\notin W$ and $c^{-}\in W$, if we swap $c^{+}$ and $c^{-}$, i.e.
-to remove $c^{-}$ from the committee and select $c^{+}$ instead, the score is increased by $\Delta(W,c^{+},c^{-})=s_{\text{PAV}}(W\setminus\{c^{-}\}\cup\{c^{+}\})-s_{\text{PAV}}(W).$
-Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
-Output: A winning committee $W$ of size $k$
-1 $W\leftarrow k$ arbitrary candidates from $C$
-2while there exist $c^{+}\notin W$ and $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$ do
-3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
-return $W$
-Algorithm 3 $\lambda$-LS-PAV
-Next, we prove the following observation in a similar spirit to Proposition 2, which shows that the local search variant of PAV can achieve maximum EJR degree if $n$ is sufficiently large and $\lambda$ is sufficiently small.
+For MDEJR, we prove the following observation in a similar spirit to Proposition 4, which shows that the local search variant of PAV (Algorithm 2) can achieve maximum EJR degree if $n$ is sufficiently large and $\lambda$ is sufficiently small.
 Proposition 5.
-Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$.
-For all $\lambda<\frac{n}{k^{2}}$ satisfying $n\geq k(k+1)(c_{\emph{max}}-1)+\lambda k^{2}$, $\lambda$-LS-PAV will output the committee achieving EJR degree $c_{\emph{max}}$.
+Given any instance with the maximum achievable EJR degree of $c_{\max}$.
+For any $\lambda\in[0,\frac{n}{k^{2}})$ satisfying $n>k(k+1)(c_{\max}-1)+\lambda k^{2}$, $\lambda$-LS-PAV will output the committee achieving EJR degree $c_{\max}$.
 Proof.
-We prove it by contradiction.
-Suppose the committee, $W$, output by $\lambda$-LS-PAV provides the EJR degree strictly smaller than $c_{\text{max}}$.
-There exists a $\ell$-cohesive group $V\subseteq N$, such that less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, i.e., |$|\{i\in V:|A_{i}\cap W|\geq\ell\}|<c_{\text{max}}$.
-Since $V$ is $\ell$-cohesive, there exist $\ell$ candidates approved by all voters in $V$.
-At least one such candidate, $c^{+}\in\bigcap_{i\in V}A_{i}$, is not selected, as otherwise all voters in $V$ approve at least $\ell$ members of $W$.
-We will show that there exists a candidate $c^{-}\in W$ that the increment of the score by swapping $c^{+}$ and $c^{-}$ is at least $\lambda$, so $W$ cannot be an output of $\lambda$-LS-PAV (since the $while$ loop can be executed continually), a contradiction.
-To see this, we try to swap $c^{+}$ and any candidate $c^{-}\in W$.
-Since $c^{+}\in A_{i}$ for all $i\in V$, we have
-$\Delta(W,c^{+},c^{-})$
-$=$ $\sum_{\begin{subarray}{c}i:c^{+}\in A_{i}\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{\begin{subarray}{c}i:c^{+}\notin A_{i}\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
-$\geq$ $\sum_{i\in V:c^{-}\notin A_{i}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V:c^{-}\in A_{i}}\frac{1}{|A_{i}\cap W|},$
-and, by summing up $\Delta(W,c^{+},c^{-})$ for $c^{-}\in W$,
-$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$
-$\geq$ $\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in V:\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in N\setminus V:\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
-$=$ $\sum_{i\in V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{i\in N\setminus V}\sum_{c^{-}:\begin{subarray}{c}c^{-}\in W\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
-$=$ $\sum_{i\in V}\frac{k-|A_{i}\cap W|}{|A_{i}\cap W|+1}-n+|V|=\sum_{i\in V}\frac{k+1}{|A_{i}\cap W|+1}-n$
-$\geq$ $\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{|A_{i}\cap W|+1}-n\geq\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\!\frac{k+1}{\ell}-n.$
-Since less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, the number of voters in $V$ approves less than $\ell$ members of $W$ is at least $|V|-c_{\text{max}}+1$. Thus,
-$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$ $\geq\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}-n$
-$\geq(k+1)\frac{|V|-c_{\text{max}}+1}{\ell}-n$
-$\geq(k+1)(\frac{n}{k}-c_{\text{max}}+1)-n$
-$=\frac{n}{k}-(k+1)(c_{\text{max}}-1)\geq\lambda k.$
-From the pigeonhole principle it follows that there exists a candidate $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$. ∎
-Since Proposition 3 holds for every initial committee $W$ in $\lambda$-LS-PAV.
-By considering the initial committee $W$ being the one with the maximum PAV-score and $\lambda=0$ (in which case the while-loop is never executed), we have the following corollary.
+$n>k(k+1)(c_{\max}-1)+\lambda k^{2}$ implies $c_{\max}<\frac{n}{k(k+1)}-\lambda\frac{k}{k+1}+1$. Proposition 3 implies this proposition. ∎
+Since Proposition 5 holds for every initial committee $W$ in $\lambda$-LS-PAV, by considering the initial committee $W$ being the one with the maximum PAV-score and $\lambda=0$ (in which case the while-loop is never executed), we have the following corollary.
 In contrast to our counterexample in Appendix C, this corollary shows the success of PAV for large enough $n$, which may be of independent interest.
 Corollary 1.
-Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$. If $n\geq k(k+1)(c_{\emph{max}}-1)$, PAV has an EJR degree of $c_{\emph{max}}$.
+Given any instance with the maximum achievable EJR degree of $c_{\max}$. If $n>k(k+1)(c_{\max}-1)$, PAV has an EJR degree of $c_{\max}$.
 By setting $\lambda=\frac{n}{k(k+1)}$ in Proposition 5, we have the following corollary, which is crucial for our algorithm.
 Corollary 2.
-Given any instance $(N,C,\mathbf{A},k)$ with maximum achievable EJR degree $c_{\emph{max}}$.
-If $n\geq k(k+1)^{2}(c_{\emph{max}}-1)$, $\frac{n}{k(k+1)}$-LS-PAV outputs a committee with EJR degree $c_{\emph{max}}$.
+Given any instance $(N,C,\mathbf{A},k)$ with maximum achievable EJR degree $c_{\max}$. If $n>k(k+1)^{2}(c_{\max}-1)$, $\frac{n}{k(k+1)}$-LS-PAV outputs a committee with EJR degree $c_{\max}$.
 Based on Corollary 2, we can use a similar way to MDJR to design our algorithm.
-In particular, we can decide whether there exists a committee providing EJR degree $c$ by enumerating all the committees when $n\leq k(k+1)(c_{\text{max}}-1)$, which can be computed in running time $f(k,c)$.
-Our algorithm is presented in Algorithm 4, which achieves the maximum EJR degree and runs in time $f(k,c_{\text{max}})\cdot\text{poly}(m,n)$.
+In particular, we can decide whether there exists a committee providing EJR degree $c$ by enumerating all the committees when $n\leq k(k+1)^{2}(c_{\max}-1)$, which can be computed in running time $f(k,c)$.
+Our algorithm is presented in Algorithm 4, which achieves the maximum EJR degree and runs in time $f(k,c_{\max})\cdot\text{poly}(m,n)$.
 The correctness of our algorithm follows from arguments similar to MDJR.
@@
 1 $W\leftarrow\frac{n}{k(k+1)}\text{-LS-PAV}(\mathcal{I})$
-2for $c:\lceil\frac{n}{k(k+1)}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
+2for $c:\lceil\frac{n}{k(k+1)^{2}}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
 3       Enumerate all those $\binom{m}{k}$ possible committees to see if there is a committee $W^{*}$ achieving EJR degree $c$;
@@
 1.3.
-We believe both the “number version” and the “ratio version” are worth-studied.
+We believe both the “number version” and the “ratio version” are worth studying.
 Which choice is better depends on the specific applications.
+Acknowledgments
+The research of Biaoshuai Tao was supported by the National Natural Science Foundation of China (No.
+62472271) and the Key Laboratory of Interdisciplinary Research of Computation and Economics (Shanghai University of Finance and Economics), Ministry of Education.
+The authors gratefully appreciate the valuable comments from IJCAI’24, AAAI’25, and AAMAS’25 reviewers, which are very helpful in improving this paper.
+We especially appreciate that one of the AAMAS’25 reviewers points out the $\frac{1}{k}$-approximation of the greedy algorithm, which leads to Theorem 1 and 2 in this paper.
 Appendix A Parameterized Complexity Basics